### v2

# Understanding Differences in News Article Interaction Patterns on Facebook: Public vs. Private Sharing with Varying Bias and Reliability

Abstract
The rapid growth of news dissemination and user engagement on social media has raised concerns about the influence and societal impact of biased and unreliable information. As a response to these concerns, a substantial body of research has been dedicated to understanding how users interact with different news. However, this research has primarily analyzed publicly shared posts. With a significant portion of engagement taking place within Facebook’s private sphere, it is therefore important to also consider the private posts. In this paper, we present the first comprehensive comparison of the interaction patterns and depth of engagement between public and private posts of different types of news content shared on Facebook. To compare these patterns, we gathered and analyzed two complementary datasets: the first includes interaction data for all Facebook posts (private + public) referencing a manually labeled collection of over 19K news articles, while the second contains only interaction data for public posts tracked by CrowdTangle. As part of our methodology, we introduce several carefully designed data processing steps that address some critical aspects missed by prior works but that (through our iterative discussions and feedback with the CrowdTangle team) emerged as important to ensure fairness for this type of study. Our findings highlight significant disparities in interaction patterns across various news classes and spheres. For example, our statistical analysis demonstrates that users engage significantly more deeply with news in the private sphere compared to the public one, underscoring the pivotal role of considering both the public and private spheres of Facebook in future research. Beyond its scholarly impact, the findings of this study can benefit a wide range of stakeholders, including Facebook content moderators, regulators, and policymakers, contributing to a healthier online discourse.
keywords:
News sharing, Facebook, User engagement, Bias, Reliability, Public vs private
1 Introduction
Social media has become a dominant medium for news consumption and sharing, with Facebook being the most popular and widely used platform for these purposes [1, 2]. For example, according to a 2021 Pew Research Center survey [2], 66% of all Americans regularly use Facebook, and almost half of them (31% of the US population) regularly get their news via the platform. Other popular platforms that users regularly use for news consumption include YouTube (22% of the US population), Twitter (13%), and Instagram (11%). This significant reach highlights the potential impact that news consumption and sharing on Facebook (and social media in general) can have on the opinions and thoughts of both individuals and the society [3].
Unfortunately, not all news are reliable. With an increasing amount of misinformation being circulated on Facebook, it is crucial to understand how users engage with news of varying reliability [4]. Improved insights and understanding into these dynamics are expected to benefit media researchers, journalists, content moderators, and policymakers, which choices based on these insights, in turn, are expected to influence public opinion and the behavior of regular users [5, 6].
Previous research has extensively examined public user engagement dynamics with different news content and the factors affecting this engagement  [7, 8, 9, 10, 11, 12]. However, with a growing privacy inclination of Facebook users and a rise in user engagement in private spaces of Facebook [13, 14], it is becoming increasingly important to understand these dynamics also with regards to private sharing. Yet, the current literature predominantly focuses on public posts, leaving gaps in our understanding of engagement differences between these realms. Most importantly, prior research has not studied the variations in public vs. non-public interaction with regards to (1) reliable and unreliable news content, (2) news with different political biases, or (3) the depth of the interactions associated with each of these types of posts and content types. These are the primary gaps addressed in this paper.
In this paper, we examine the differences in public and private sharing of news articles on Facebook, with an emphasis on users’ interactions with these articles. Specifically, we investigate how news articles written with different bias and reliability are shared and engaged with, as well as the depth of these interactions. In this context, bias refers to a tendency for news articles to exhibit partiality or favoritism towards particular groups or ideas (e.g., left or right on the political spectrum), while reliability refers to the accuracy and credibility of the information presented [15] (e.g., fake or true). While most previous research focuses on either bias or reliability, we consider both dimensions, as they have been found related, but yet each has its own ability to influence news article sharing behaviors and affect the quality and diversity of information that users encounter on social media [16, 17]. On the other hand, by considering whether an interaction is deep or shallow, we can identify whether there are areas (private or public) where users engage more deeply with specific content types (e.g., fake news). This is important since the depth of user engagement can significantly influence the success of an article. For example, an article may have greater success if people share and comment on the posting of the article than if they only press Facebook’s “like” button [1].
Research questions and methodology: With the above goals in mind, we designed our study to address the following research questions:
  1. RQ1
    How do the patterns of public and private sharing of news articles on Facebook differ for articles with varying levels of bias and reliability?
  2. RQ2
    If and how does the depth of user interactions with news articles differ within public and private sharing contexts for articles with varying levels of bias and reliability?
We took several steps to address these previously unaddressed questions in as controlled matter as possible. First, in contrast to most prior works that use publisher-level labeling of news (e.g., Adelson et al. [7]; Horne et al. [18]), we employ article-level labeling of bias and reliability. By doing so, we capture that not all articles by a publisher have the same bias or reliability, recognizing that the publisher (source) is only one factor [19] in the bias/reliability. In particular, we consider the bias and reliability of articles ($N=19,050$), meticulously selected from a substantial pool of publishers ($N=1,121$) and 30K+ manually labeled news articles. Here, the original labeling was provided by Ad Fontes Media [20], which rates both news sources and individual articles based on their level of bias and reliability. Second, through active back-and-forth discussions and feedback with the Crowdtangle team, we developed and implemented a methodology using their Chrome addon that allows us to obtain simultaneous interaction statistics for a representative sample of publicly shared Facebook posts linking these articles, as well as across all Facebook posts (including both public and private posts) linking these articles. Finally, using this unique dataset, we perform a comparison of interaction dynamics when users share news articles with different levels of bias and reliability, both publicly and privately.
Empirical example findings: Our analysis reveals several key insights into how users interact with news articles of varying bias and reliability on Facebook. For instance, we show that users tend to engage more deeply in private discussions than in public ones, irrespective of the news class. When considering the news class, we highlight that users exhibit relatively higher deep interaction levels with highly-unreliable content. Our results also show that reliable news content has significantly lower private interaction shares compared to the highly-reliable or even unreliable content.
Example beneficiaries: Our methodology and findings contribute to a better understanding of the dynamics of news sharing and interaction on Facebook and provide valuable insights for multiple stakeholders. For example, considering the significant impact that social media interactions have on content and platform popularity [21], our findings on how bias and reliability influence user engagement in public vs. private spheres can aid media researchers and journalists. These insights can inform more effective news content creation and distribution strategies for enhanced audience engagement. For content moderators of platforms such as Facebook, the analysis offers data-driven guidance on where moderation efforts may be most needed to curb the spread of problematic content. For policymakers and regulators, our findings illustrate how user privacy settings impact news content engagement types and interaction depth. Notably, highly unreliable news receives more private discussion engagement share compared to the broader unreliable category. These results underscore how platform design choices can influence selective exposure and engagement, potentially exacerbating issues like ideological polarization and misinformation spread.
Roadmap: Sect. 2 explains our multi-faceted research design, including news article selection, labeling, interactions data collection, and the processing steps taken to ensure fair comparisons between public and private engagement. The resulting dataset is summarized in Sect. 3. Sect. 4 analyzes public vs. private dynamics both from a high-level aggregated perspective (Sect. 4.1) and then using a detailed statistical analysis (Sect. 4.2). In Sect. 5, we turn our attention to differences in the depth of interactions. Sect. 6 concludes our analysis with a correlation-based analysis between public and private interactions. Discussion of the key findings is provided in Sect. 7. Finally, related works (Sect. 8) and conclusions (Sect. 9) are presented.
2 Research Design, Data Collection and Limitations
At a high level, our research design has three parts.
  • 1.
    Article selection and labeling: We first obtain the URLs to a substantial number of news articles, which have been classified by Ad Fontes Media [20] with regards to both their individual bias and reliability. We then decide category thresholds and label each article into one of three bias classes and one of four reliability classes.
  • 2.
    Collection of interactive data: After careful preprocessing of the URLs, we use the CrowdTangle browser extension to obtain interaction data for two sets of posts linked to these URLs (1) public and (2) combined (public + private), where the first set is a subset of the second. To address some limitations of the API and to ensure fair comparison of the sets, we leave a four-month gap between the latest labeled article (Nov. 1, 2022) and the primary data collection (Mar. 2023), as well as apply some additional post-processing (e.g., examining the actual posts), filtering (e.g., based on thresholds determined via communication with the CrowdTangle team), and collect some complementing data directly from CrowdTangle (e.g., to address limitations of the extension API).
  • 3.
    Data analysis: Finally, we use the final dataset (capturing the two sets) to compare the properties of the public vs. combined (private + public) sets and discover statistically significant differences between the interaction patterns of public and private posts. Here, it should also be noted that the combined set is dominated by private posts.
We next provide detailed descriptions of the initial two steps, before presenting our analysis results (step 3) in the subsequent sections.
2.1 News Article Selection and Labeling
Selection of articles: There are several independent initiatives that assess the bias and/or reliability of individual news articles and/or news sources. Examples of such evaluation efforts include Media Bias Fact Check [22], Ad Fontes Media, AllSides [23], and NewsGuard [24]. Among these, we opted to use data from Ad Fontes Media for the following main reasons: (1) it has been widely used in previous research [25, 26, 27, 12], (2) they evaluate individual news articles (not only the news publishers), (3) each assessed article receives a score for both bias and reliability, (4) they offer a transparent evaluation methodology, which is published and explained in a white paper [28], and finally (5) the dataset comprises a large set of news contents (including news articles, podcasts and TV shows) randomly selected from a diverse set of publishers.
We obtained all evaluated news from Ad Fontes Media as of Nov. 1, 2022. After filtering the podcasts and TV shows, this set included 31,446 news article URLs from 1,121 publishers. We also filtered a few URLs that (1) primarily served as event reporting pages (e.g., https://www.nola.com/news/hurricane) and (2) articles for which Ad Fontes Media had changed their rating (identified through complementary documents published by Ad Fontes Media). After excluding these articles, the dataset comprised 31,408 articles.
Labeling of articles: Each article in the Ad Fontes Media dataset is assessed for both bias and reliability by a minimum of three human analysts, representing a mix of right, left, and center self-reported political perspectives. Bias scores from Ad Fontes Media range between -42 and +42, where more negative values suggest a stronger left-ward bias, and positive values indicate a right-leaning bias. As for reliability, scores vary from 0 to 64, with 64 representing the most-reliable news.
Figure 1: CDF of bias scores
Figure 2: CDF of reliability scores
For our analysis, we categorized the bias and reliability scores into distinct classes. While previous work mainly focuses on binary classification of news (e.g., Left vs. Right or Fake vs. True), we opted for a more granular approach by defining three classes of bias (Left, Right, and Center) and four classes of reliability (Most-unreliable, Unreliable, Reliable, and Most-reliable). Figures 1 and 2 show the Cumulative Distribution Function (CDF) of the bias and reliability scores of the articles included in our analysis (after additional filtering steps, explained later in this section) together with the threshold values and class labels used to define each class.
With our labeling, anything with a bias score below -6 is called Left-biased, and anything above 6 is called Right-biased, with the range (-6,6) capturing the Center class. We note that Ad Fontes Media also refers to this range as “Middle or Balanced Bias” [20]. As observed, this class accounts for most samples in our dataset, representing 58% (11,094 out of 19,050 articles in the final dataset). With this split, the two biased classes have similar, close to 20%, shares.
For the reliability classes, we assigned the Most-unreliable class to cover the lowest range (0, 32). According to Ad Fontes Media’s terminology [20], these values typically correspond to articles “which contain inaccurate and misleading info, selective stories or have opinion and wide variation in their reliability”. Similarly, the range (48-64), chosen for the Most-reliable class, encompasses articles characterized by “thorough and original fact reporting” due to Ad Fontes terminology. Finally, the threshold of 40, between Reliable and Unreliable, matches the mid-point between these two classes as well as where Ad Fontes make their split between “wide variation in reliability” and “mix of fact reporting and analysis”.
2.2 Collection of Interaction Data
Preprocessing of URLs: Before collecting the interaction data associated with each article (URL), we (1) replaced the link shorteners with their corresponding URLs and (2) converted all article URLs to their canonical forms. During the conversion process, we primarily removed all unnecessary URL parameters (those sent after the ”?” symbol) except for the ones essential to the canonical form (e.g., “id” parameters are sometimes part of the canonical form). As another step in identifying the canonical forms, we replaced all the web archive links in the Ad Fontes dataset with the original links.
Primary data collection for the two sets of posts: We used the CrowdTangle browser extension [29] to collect the interaction data for the Facebook posts sharing each news article of interest. In addition to interaction statistics for the public posts that CrowdTangle (owned by Facebook) indexes and tracks over time (their database, available for and used by researchers, contains longitudinal interaction data for indexed posts), this extension provides access to the life-time interaction statistics, obtained thus far, across all posts (private + public). In this paper, we use this extension to collect statistics about all interactions associated with these two sets of posts: combined and public, where the combined captures all interactions of all posts (public + private) and the public is a subset of the combined (only capturing public posts).
To collect the data for each URL, we used the “Download” option within the extension’s interface when browsing each article. This allowed us to easily download a CSV file containing the interactions data for both sets. As expected from a privacy standpoint, only the aggregated statistics are provided for the combined set. In contrast, every post and their individual interaction data are provided for the public set.
At a high level, the combined category aggregates the interactions (e.g., likes, shares, comments, etc.) of all Facebook posts referencing the article’s URL, regardless of the post’s privacy settings. This includes interactions on posts with limited privacy settings, such as “Friends only” or “Only Me”. In contrast, the public category, includes all interactions with the public posts tracked by CrowdTangle, which, due to heavy-tailed characteristics and significant coverage, captures most interactions with public Facebook posts. As supporting examples, all posts of all US-based public groups with 2K+ members are indexed, and in 2021 over 99% of all Facebook pages with at least 25K likes were indexed [30]. While some public posts are not indexed, we note that their indexing does not consider any bias/reliability criteria, is expected to capture the majority of public posts, and the majority of the posts in the combined set are private.
Timeline and limitations: We collected the interaction data in Mar. 2023, ensuring at least a four-month gap between the publication date of the newest article in our dataset and our interaction data collection. This length of interval was chosen to ensure that we capture most of the posts and interactions for the URLs in our dataset.
As mentioned in [31] and confirmed via our communication with CrowdTangle, the information from the CrowdTangle addon originates from two different sources: (1) public interaction data is obtained from the CrowdTangle database, while (2) combined interaction data is derived from Facebook’s Graph API. Consequently, variations in data capturing times might cause some URLs to exhibit higher public than combined interactions [31]. Furthermore, as discussed in [31], Facebook’s aggregation data restarts the interaction count whenever websites update their connection schemes (e.g., switching from HTTP to HTTPS) after publishing a URL, while CrowdTangle continues accumulating data. Although the discrepancy is not substantial (9% in our initial dataset), we aimed to mitigate the effect of these differences where possible (described next).
Further filtering and data processing for fairer comparisons: After consulting with the CrowdTangle team regarding all numerical discrepancies that we observed, we identified two additional reasons for the discrepancies. First, as stated in [32], for privacy reasons, for the combined data (that include private posts) “the values are intentionally not precise, but you can be confident they accurately reflect user engagement with a URL.” Discussing the identified discrepancy for such cases, the CrowdTangle team confirmed that combined interaction data with counts below 100 is not reliable; however, that data with counts above 100 provides a reliable estimate of public + private interactions. For this reason, we removed all URLs in our dataset with combined interactions below 100. Although this reduced our dataset to 19,505 articles, it is important to note that this group of URLs accounted for less than 4% of public interactions in our dataset. Furthermore, we note that the public interaction data for this group still remains reliable.
Second, the CrowdTangle team informed us that the Facebook Graph API only considers posts with the URL attached to the post (i.e., those displaying a preview of the URL) when counting interaction with the URL, while CrowdTangle’s algorithm includes any post containing a linked URL, regardless of attachment status. Taking this into account, in our final step toward enhancing interaction data quality, we chose not to rely on the aggregated data from the CrowdTangle extension (which sums the interactions of all posts mentioned in the extension). Instead, we examined each post to determine whether the related URL was attached or not, and if so, we included them in the sum of public interactions for that URL. [^1] The attachment status of all the posts and URLs will be published with the camera-ready version for the interested researchers. This resulted in a more accurate comparison between combined and public interactions. After these processing steps, we had only 455 URLs for which public interactions exceeded combined interactions, which were excluded from our final dataset. The primary causes of these discrepancies were beyond the scope of our research to address (e.g., changes in URL, publisher protocol schemes, or different data capturing times as mentioned above).
While some of the pre- and post-processing steps described above require significant effort, they help ensure an accurate and fair comparison between the public and combined sets in such a way that we can provide conclusive insights into the relative sharing patterns of public vs. private posts.
Enhancing the data with additional posts: Another limitation of the CrowdTangle extension is that it only allows us to retrieve data for up to 500 posts per URL. To address this limitation, for the URLs with more than 500 posts in the public interaction category, we obtained additional data directly from CrowdTangle. However, these URLs constitute only 42 instances in our dataset and will have no significant impact on the reproducibility of our statistical analysis.
3 Dataset
High-level summary: After applying the aforementioned filtering steps, we have in total 19,050 articles from 1,121 news outlets remaining in our dataset. These articles have been shared in 253,350 posts, which combined are responsible for 225,282,745 interactions (out of which 70,428,316 are public). Table 1 summarizes the dataset.
Bias data: Table 2 shows the number of articles and the total combined and public interactions for each bias and reliability class. With the selected bias thresholds, our dataset includes 4,516 Left-biased articles and 3,440 Right-biased articles, with the remaining 11,094 articles falling in the Center category. These articles are in turn responsible for 79,038,246 (Left), 56,160,101 (Right), and 90,084,398 (Center) interactions. While the Center category includes 58% of the articles, it is interesting to note that it is responsible for a significantly smaller fraction of the total interactions (40%). Instead, the Left-biased and Right-biased articles see relatively higher interaction rates (studied in the next section). For example, the Left-biased articles account for only 24% of the articles but 35% of the interactions, and the Right-biased articles are responsible for only 18% of the articles but 25% of the interactions.
Table 1: Summary of the final dataset
Type Articles Publishers Posts Combined interactions Public interactions
Number 19,050 1,121 253,350 225,282,745 70,428,316
Table 2: Statistics based on Bias and Reliability
Class Articles no Combined interactions Public interactions
Bias Left 4,516 79,038,246 20,307,107
Center 11,094 90,084,398 32,109,924
Right 3,440 56,160,101 18,011,285
Reliability Most-unreliable 3,520 45,299,475 13,206,643
Unreliable 3,795 46,712,927 14,199,322
Reliable 9,308 89,596,137 30,639,973
Most-reliable 2,427 43,674,206 12,382,378
Reliability data: For the reliability classes as shown in Table 2, the relative differences are smaller. Here, the Reliable articles make up the largest share (9,308 articles and 89,596,137 interactions), followed by Unreliable (3,795 articles and 46,712,927 interactions), Most-unreliable (3,520 articles and 45,299,475 interactions), and Most-reliable (2,427 articles and 43,674,206 interactions).
4 Public vs. Private Interactions
To better understand the differences in interactions with public vs. private posts, we first compare the interaction rates with all public posts and with all combined (public + private) posts. We note that this implicit comparison approach provides more stable and reliable statistical comparisons than trying to use only the estimated number of private interactions. The relative strength of this approach is further compounded by the private interactions typically dominating the combined set.
4.1 Aggregated (Macro) Analysis of Interactions
Figure 3: Normalized number of combined and public interactions per article for each bias and reliability class.
Figure 3 presents the number of combined and public interactions per article for each class. Here, two metrics are used:
  • Average combined interactions (left bars for each class) represents the average number of combined (private+public) interactions per article in each class. More specifically, considering $I_{c,i}^{\text{comb}}$ as the combined interaction for the $i^{th}$ article in class $c$, the average combined interactions for this class ($\bar{I}_{c}^{\text{comb}}$) is calculated as:
    $\bar{I}_{c}^{\text{comb}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}{N_{c}},$ (1)
    where the numerator ($\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}$) corresponds to the “Combined interactions” column in Table 2 and the denominator (i.e., the total number of articles $N_{c}$ for class $c$) is found in the “Articles no” column of Table 2. For example, for the Center class in the Bias category, the average combined interactions per article is computed by dividing the total interactions (90,084,398) by the articles count (11,094), yielding $\bar{I}_{\text{Center}}^{\text{comb}}$ to be 8,120 combined interactions per article (gray bar in Figure 3).
  • Average public interactions per article (right bars) is defined in a similar manner but considering only the public interactions. More specifically, considering $I_{c,i}^{\text{public}}$ as the public interactions for the $i^{th}$ article in a class $c$, the average public interactions for this class ($\bar{I}_{c}^{\text{public}}$) is calculated as follows:
    $\bar{I}_{c}^{\text{public}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{N_{c}},$ (2)
    where the numerator (i.e., $\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}$) corresponds to the “Public Interactions” column in Table 2. Now, taking the Center class as an example again, the average public interactions per article is calculated by dividing the total public interactions (32,109,924) by the article count (11,094), resulting in an $\bar{I}_{\text{Center}}^{\text{public}}$ of 2,894 interactions per article (striped gray bar in Figure 3).
As seen in the figure, there are some very clear and interesting differences in the trends observed for the public vs. the combined statistics. First, for all three bias classes and for all four reliability classes, the interaction rates are significantly lower for the public posts than for the combined set (which for all classes are dominated by private posts). These aggregate rate differences suggest that users are more likely to interact with private posts, regardless of bias and reliability class, and underscores the importance of considering the total (combined) interactions, not only the (typically studied) public posts. We next look closer at the relative interaction levels of each class and the impact of using the public vs. combined sets for such interaction comparisons.
Bias comparisons: Second, looking at the relative differences among the bias classes, the Left bias class has the highest average combined interactions per article (17,501), followed by the Right (16,325) and Center (8,120) classes. For public interactions; however, the Right bias class has the highest average public interactions per article (5,236), with the Left (4,497) and Center (2,894) classes trailing behind. This shows that biased articles see even greater relative interaction rates when posts are private than public.
Reliability comparisons: Third, for the reliability classes, the Most-reliable class has the highest average combined and public interactions per article (17,995 and 5,101). The Unreliable and Reliable classes have relatively similar average public interactions per article (3,741 and 3,291, respectively). Moreover, the Most-unreliable class has a higher average combined interactions per article (12,869) and public interactions per article (3,751). It is interesting to see that the two extreme classes (i.e., the Most-reliable and Most-unreliable) have the highest average interactions per article in both combined and public contexts, implying that users are more likely to engage with articles on both ends of the reliability spectrum, regardless if the posts are public or non-public, although the level of interactions is different.
Key observation: Compared to the Center group, biased groups receive higher levels of aggregated engagement per article. In terms of reliability, the Reliable news reaches the lowest engagement levels per article.
Aggregated public interaction share: Motivated by most prior works only studying the public posts tracked by CrowdTangle, we next look closer at how big fraction of the total interactions this set captures and compare the relative differences in the public interaction share of the different categories. For this analysis, we calculate the (aggregate) public interaction share, defined as the ratio of public interactions to the combined interactions for each class $c$ as $\bar{I}_{c}^{\text{public}}/\bar{I}_{c}^{\text{comb}}$.
Figure 4: Aggregated public interactions shares
Figure 4 summarizes these ratios. We note that the Center class has the highest aggregated public interaction share, and that among the biased groups, there is a clear gap between the left and right parties, as the Right class has a significantly (25% extra) higher public interaction share (lower private interaction share). When it comes to the reliability classes, the relative gap between the classes is smaller. However, it should be noted that the two extreme classes (Most-reliable and Most-unreliable) have the lowest (and almost similar) aggregated public interaction shares.
Key observation: Compared to right-biased articles, left-biased articles receive higher private interactions share. In terms of reliability, extremely reliable/unreliable news articles receive higher levels of private interactions share.
4.2 Granular Statistical (Micro) Analysis of Interactions
Having presented aggregated analysis and insights, we next present a more detailed statistical comparison between different classes in which we give equal consideration to every article in a class, regardless of its total interaction count. For this analysis, we utilize the distributions of the per-article level interactions for each class $c$: $\mathcal{D}^{\text{comb}}_{c}=\{I_{c,i}^{\text{comb}}\}$ and $\mathcal{D}^{\text{public}}_{c}=\{I_{c,i}^{\text{public}}\}$, where $I_{c,i}^{\text{comb}}$ is the total number of interactions across all posts associated with the $i^{th}$ article of class $c$ ($1\leq i\leq N_{c}$) and $I_{c,i}^{\text{public}}$ is the corresponding number of interactions with public posts.
The next two subsections compare the distributions of the total (combined) interactions (Sect. 4.2.1) and study the distributions of the relative public interaction shares (Sect. 4.2.2), respectively. In both subsections, the distributions are compared using multiple statistical tests; statistical significance is reported if a p-value is below 0.001, and supporting p-values are reported for the main findings.
4.2.1 Distribution of Combined Interactions
[…]
Figure 5: Empirical complementary cumulative distribution function (CCDF) of both combined and public interactions, alongside an inset boxplot detailing the distribution of the combined interactions for different bias classes. While the CCDF underscores the tail characteristics of the distributions, the boxplot provides a granular statistical summary of the combined interactions, including medians, quartiles, and means.
Upon examining the figure, we can make several observations. First, as articles with fewer than 100 combined interactions were excluded, the minimum value for $\mathcal{D}^{\text{comb}}_{c}$ starts at 100. Second, the shape of the of the CCDFs when plotted on log-log scale underscores the “heavy-tailed” nature of the interactions ($\mathcal{D}^{\text{comb}}_{c}$ and $\mathcal{D}^{\text{public}}_{c}$) of all bias classes $c$. This suggests that a small subset of the articles are responsible for a most of the interactions. Third, there are discernible variations in the distributions. For example, articles classified as biased (both Left and Right) consistently see higher interaction levels (i.e., right-shifted curves compared to the Center class), echoing the trends we observed in the aggregated analysis. To better understand these disparities and their implications, we next provide a more rigorous statistical analysis.
Statistical Tests for Bias-related Differences: To evaluate the differences in distributions of the set of combined interactions $\mathcal{D}^{\text{comb.}}_{c}$ across all bias classes $c$, we employed a multi-step statistical testing approach:
  1. Overall Differences: We first used the Kruskal-Wallis test to identify overall differences between the distributions between the bias classes. This test revealed significant disparities among the classes, with a p-value of $1.66\cdot 10^{-65}$.
  2. Pairwise Comparisons: Second, we used the Dunn test to identify specific pairs of classes that exhibited differences. This test confirmed that the two biased classes Left and Right are different than the Center class, but the difference between the two biased classes themselves was not significant. Here, the most significant difference were observed between the Left and Center class (i.e., $\mathcal{D}^{\text{comb}}_{\text{Left}}$ and $\mathcal{D}^{\text{comb}}_{\text{Center}}$), registering a p-value of $2.06\cdot 10^{-55}$.
  3. Median Comparisons: Third, we used the Mann-Whitney U test to compare the medians of the three classes: Left (1,440), Center (842), and Right (1,286). Also, here, we observed statistically significant pairwise differences when comparing each of the two biased classes (Left and Right) with the Center class, but not between each other. Here, the two significant cases obtained p-values of $8.4\cdot 10^{-57}$ and $9.1\cdot 10^{-28}$, respectively.
  4. Comparing Means: Finally, due to the violations of assumptions intrinsic to a t-test (namely, normality and homogeneity of variances), we instead use bootstrapping with 100K iterations and a 99% confidence interval to compare the means. Again, the pairwise differences in the means between Left vs. Center and between Right vs. Center are statistically significant, but not between the two biased classes themselves (i.e., Left vs. Right).
Key observation: Statistical analysis further strengthens the observation that biased news consistently garners higher combined engagement levels per article than center-aligned news. Interestingly, there is no statistically significant difference between the Left and Right class.
Comparison of Reliability Classes: Figure 6 shows the CCDFs for the different reliability classes, broken down for both combined and public interactions, accompanied by a boxplot offering a clearer perspective on the percentile values for the combined interactions. The CCDFs again demonstrate heavy-tailed distribution characteristic, and the relative shifts of the distributions are consistent with the aggregate (average values) previously discussed (and seen in Figure 3), with the Most-reliable class typically getting the most interactions, followed by the unreliable classes, and subsequently by the Reliable class. While comparing the tail parts is straightforward, comparing the whole distributions requires some care. For example, here, the core of the boxplot does not distinctly set apart the Most-reliable from the unreliable classes. To derive robust conclusions, we therefore again performed a sequence of statistical tests on the combined interactions distributions.
Figure 6: CCDF of both combined and public interactions, alongside an inset boxplot detailing the distribution of the combined interactions for different reliability classes. While the CCDF underscores the tail characteristics of the distributions, the boxplot provides a granular statistical summary of the combined interactions, including medians, quartiles, and means.
First, the Kruskal-Wallis test was applied to discern differences among the reliability classes. While we observed statistical differences (p-value of $7.9\cdot 10^{-20}$), the magnitude of distinction among reliability classes is more nuanced than in the bias classes (which had a p-value of $1.66\cdot 10^{-65}$). Further analysis employing the Dunn’s test confirmed that the Reliable class statistically diverges from the other three. The least significant p-value here, $7.63\cdot 10^{-8}$, is attributed to the comparison between the Reliable and Unreliable classes. The most significant, on the other hand, emerges from the Reliable versus Most-reliable comparison with the p-value of $7.05\cdot 10^{-15}$. Except for these distributions, none of the other pairwise distribution comparisons resulted in statistically significant differences. Third, the Mann-Whitney U test reinforced our findings, particularly highlighting the statistical significance when comparing the medians of the Reliable class (median of 891) with the others: Most-unreliable (1,162), Unreliable (1,104), and Most-reliable (1,218). While all three pairwise comparisons are significant, their p-values range from $8.96\cdot 10^{-9}$ (Reliable vs. Unreliable) to $8.4\cdot 10^{-16}$ (Reliable vs. Most-reliable). Finally, we apply bootstrapping to compare the means. While only the Reliable and Most-reliable classes comparison is significant at the 99% confidence level, we note that all three comparisons against the Reliable class are significant at the 95% confidence level.
Key observation: Comparing to the other reliability classes, the Reliable articles demonstrate statistically significant lower interaction values, especially when compared to the Most-reliable articles.
4.2.2 Distribution of Public Interaction Shares
The preceding analysis centered on the distribution of combined interactions. Like the aggregated analysis, we next shift our focus to the proportion of interactions that are public. Specifically, for each class, we analyze the distribution of public interaction shares at the granularity of individual articles. For this analysis, we define the set of public interaction ratios (PIR) of class $c$ as $\mathcal{D}^{\text{PIR}}_{c}=\left\{\frac{I_{c,i}^{\text{public}}}{I_{c,i}^{\text{comb}}}\right\}$, where $1\leq i\leq N_{c}$ is the index of articles included in class $c$.
Figure 7: Distribution of public interaction ratios
[…]
Motivated by the higher significance (smaller p-value) for the reliability classes, we consider these differences first. Using the Dunn test, we find statistically significant differences for all pairwise distribution comparisons except between the Most-unreliable and Most-reliable classes. Similarly, the Mann-Whitney U test shows that all pairwise relative differences in the medians (observed in Figure 7), except for between the Most-unreliable and Most-reliable classes, are significant (with the largest p-value among the these pairwise cases being $2.82\cdot 10^{-14}$). Finally, bootstrapping confirms that all pairwise differences in the means observed in the figure are statistically significant, except for the case of the Most-unreliable vs. the Most-reliable class. It is also worth noting that our observation that the two extreme classes have lower public interaction ratios (higher private interaction ratios) aligns with our findings from the aggregated analysis.
Key observation: The two extreme classes, Most-reliable and Most-unreliable, exhibit similar private interaction ratios that are statistically higher compared to the other classes.
We now shift our attention to the public interaction ratios across the bias classes, where we make some interesting observations regarding the classes’ relative order. Specifically, we now observe the first difference in the relative order of the classes when comparing the average (and median) public shares calculated on a per-article basis (Figure 7) with the aggregated results (Figure 4). More specifically, comparing with the aggregate results, the Right class has now switched ranking with the Left class to become the class with the noticeably smallest public interaction share. Furthermore, comparing the distributions seen for the Right and Left class is statistically significant: Dunn test (p-value of $3.25\times 10^{-14}$), Mann-Whitney U test (p-value of $6.16\times 10^{-17}$) and bootstrap results being significant at 99% level.
While these results may appear surprising at first, they are due to some interesting differences in the distributions. First, note that the aggregate analysis only calculates the overall ratio $\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}$ compared to the granular analysis, where the average (for example) is calculated as $\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\frac{I_{c,i}^{\text{public}}}{I_{c,i}^{\text{comb}}}$. Second, taking a closer look at the distributions, $\mathcal{D}_{Left}^{PIR}$ and $\mathcal{D}_{Right}^{PIR}$ differ substantially. Part of this can be observed in Figure 5, where we can see that Left class demonstrates notable disparities between public and combined interactions in the tail, while the Right class exhibits relatively larger differences in the head of the distribution. This suggests that articles with higher combined number of interactions in the Right class may exhibit higher normalized public interaction shares, and articles with few combined interactions may exhibit relatively smaller public shares (compared to the Left class).
To confirm this conjecture and substantiate our claim that the above differences are due to aggregated statistics being more affected by larger public interaction shares associated with the tail, we compared the public interaction ratios for articles associated with the tail and head of the distributions of the combined number of interactions. For this analysis, we split the total sets ($\mathcal{D}^{PIR}_{Left}$ and $\mathcal{D}^{PIR}_{Right}$) into two to five equally sized subsets (based on the total combined number of interactions $I_{c,i}^{comb}$ associated with each entry) and then compared the medians and averages of the public interaction ratios seen in the first and last bucket of the Right class with the corresponding values seen for the Left class. In all cases, the Left dominates the Right for the head and is dominated by the Right for the tail. For example, with three equal-sized buckets, the two classes have median ratios of 0.34 vs. 0.13 for the first bucket (head) and 0.39 vs. 0.46 for the last bucket (tail). Furthermore, we observed statistically significant differences (99% confidence level) as per the above conjecture with both the Mann-Whitney U test (medians) and Bootstrap tests (averages), validating these differences. This confirms that the articles in the Right class with the highest interactions have elevated normalized public interaction ratios, and those in the Left class with the lowest interactions have similarly elevated ratios.
Key observation: In terms of bias, the Right class sees the smallest public interaction share among the least popular content to interact with, while the Left class sees relatively less overall public interaction share (due to its most popular articles to interact with not seeing as big public interaction share).
5 Deep vs. Shallow Interactions
In this section, we investigate the prevalence of deep vs. shallow interactions and explore their variations among different reliability and bias classes. For this analysis, we categorize all (emoji-based) reactions (i.e., likes, loves, sads, etc.) as shallow interactions while comments and shares are considered as deep interactions. This distinction, supported by the previous works [8, 33], is motivated by the idea that comments and shares typically involve more cognitive effort and engagement from users, as they require formulating thoughts, opinions, or responses in the form of comments or actively endorsing and disseminating the content by sharing. In contrast, reacting with an emoji is a more passive and less cognitively demanding form of interaction.
To compare the depth of interaction seen for subsets of public and private posts, we define the following two per-article metrics:
  • Deep Interactions Ratio for Combined (DIRC): For each article, this ratio measures the proportion of the combined interactions that are classified as deep interactions.
  • Deep Interactions Ratio for Public (DIRP): For each article, this ratio measures the proportion of the public interactions that are classified as deep interactions.
For example, consider an article with 1,000 total interactions, 500 of which are deep interactions, resulting in a DIRC of 0.5. Now, if 400 of these interactions are public, and 100 of them are deep interactions, the DIRP is 0.25 for the same article. We next analyze the distribution of DIRC and DIRP values for all articles in each class.
Comparisons of bias classes: We first compare the deep interaction patterns of the bias classes. Figure 8 shows the distributions of DIRC and DIRP for the different classes. As a baseline, we also include the distributions for the general population (rightmost box pair), considering all samples.
Figure 8: Deep interaction ratios (DIRC and DIRP) for different bias classes and a baseline.
First, referring to the figure, for each bias class, we observe substantial difference between the distributions of DIRC and DIRP. For example, for every percentile and for the means, the DRIC distributions (combined) have significantly larger values than the DIRP distributions (public). This again highlights the difference between the public and combined interactions and the importance of considering both the private and public spheres in future research. This observation leads us to the following key insight.
Key observation: Irrespective of the news class, users tend to engage more deeply in private discussions compared to public ones.
Now, and perhaps even more importantly, to better understand to what extent the relative level of interaction seen for different bias classes depends on which dataset was used we apply the same statistical testing methodology as discussed in previous sections for the two sets independently. First, the Kruskal-Wallis test provides evidence that the distributions differ significantly among the bias classes, regardless if using of the DIRC and DIRP, with p-values of $7.14\cdot 10^{-23}$ and $2.08\cdot 10^{-9}$, respectively. Comparing the p-values, we note that the differences appear more significant using the combined set (DIRC) than using only the public data (DIRP). Second, consistent with visual comparison in the figure, the post-hoc Dunn test only supports the statistically significance differences between the Right group (which has the biggest fraction of deep interactions regardless of using the combined or pubic sets) and the other two groups but not among the Center and Left group. Finally, the Mann-Whitney U test and bootstrapping support the same pattern for the median and means, with the only significant differences again being between the Right class and the other two bias classes. For the bias classes, the main results regarding the relative depth of the interactions among the classes are, therefore, consistent regardless of the dataset.
Figure 9: CDFs of DIRC (combined) and DIRP (public) for different bias classes.
To illustrate the tendency of the posts with Right-biased articles seeing a higher deep interaction share, regardless of whether publicly shared or not, we include the CDFs of DIRC (combined) and DIRP (public) for all three biases classes in Figure 9. We note that in both cases, the Right class exhibits a noticeable shift to the right, supporting the following insight.
Key observation: Facebook users tend to engage more deeply with content from the Right bias class in both public and combined contexts.
Comparison of reliability classes: We now shift our focus to the reliability classes and their deep interaction ratios. Figure 10 summarizes these results. While we again observe deeper interactions for the combined set (suggesting shallower interactions with public posts), the relative differences among the reliability classes themselves (given a dataset) are less apparent compared to those observed among the bias classes. We also observe some smaller differences in which relative differences are significant, highlighting the value of also considering the combined data (not only the public data).
Figure 10: Deep interaction ratios (DIRC and DIRP) for different reliability classes
Figure 11: CDFs of DIRC (combined) and DIRP (public) for different reliability classes.
First, starting with the combined dataset (DIRC), we observe presence of significant differences among the distributions (Kruskal-Wallis test), with the post-hoc Dunn test (distribution) and Mann-Witney U test (median) revealing that only the Most-unreliable class differs statistically from the others. This observation is further visually supported by the CDF plot in Figure 11, which clearly shows a shift to the right in the DIRC distribution of the Most-unreliable class.
Switching our focus to the DIRP distributions, containing only public interactions, we interestingly observe a slightly different pattern, as we, for that dataset, also observe statistical differences between some additional classes. First, as suggested by the DIRP CDFs in Figure 11, both extreme classes (Most-unreliable and Most-reliable) are statistically different from the other two classes (Kruskal-Wallis test followed by Mann-Whitney test, followed by the post-hoc Dunn test). Second, no statistical significance was observed between the Most-reliable and Most-unreliable classes themselves or between the Reliable and Unreliable classes. We base the following insight based on our observations across the two datasets.
Key observation: Users exhibit higher deep interaction levels with content from the Most-unreliable class.
6 Correlation-based Analysis
Given the significant differences observed in interaction rates and deep interaction levels between public and private posts, we aim to determine if these variations stem from specific articles within each bias/reliability class that receive more or fewer private interactions or if certain articles consistently exhibit higher (deep) interaction rates, regardless of public or private sharing. To investigate this, we next examine the correlations between public post statistics and their combined counterparts.
Public vs. combined interactions: Table 3 summarizes the correlation statistics (between combined and public interactions) observed for each class using two different correlation metrics: Spearman and Kendall. Given that the normality assumption required for the Pearson correlation is not satisfied for the different classes, we opted for non-parametric correlation metrics. These metrics are better suited to capture monotonic relationships, account for outliers, and take into consideration the ranks of the data points.
Table 3: Correlation between public and combined interactions. (Bold/italics indicate highest/lowest correlation.)
Class Spearman Kendall
Bias Left 0.82 0.66
Center 0.74 0.59
Right 0.74 0.58
Reliability Most-unreliable 0.73 0.58
Unreliable 0.79 0.63
Reliable 0.76 0.61
Most-reliable 0.77 0.62
While different trends are observed for reliability classes, a consistent pattern emerges for the bias classes across both correlation metrics. The Left class exhibits the strongest correlations, followed by the Center class, and finally, the Right class with the weakest correlations. This suggests that articles with a left-leaning bias tend to have more similar sharing and interaction patterns in both public and private contexts, indicating higher similarity in engagement across different articles. In contrast, articles with a right-leaning bias show more variability in their public and private interactions, which may reflect a more diverse range of opinions or reactions to the content among readers. These observations are consistent with our observations in Section 4.2.2, which showed that the Right class saw much more diverse public interaction shares depending on the level of combined interactions of the article (see our bucket-based analysis).
Table 4: Correlations for deep interaction shares: public vs combined. (Bold/italics indicate highest/lowest correlation.)
Class Spearman Kendall
Bias Left 0.64 0.48
Center 0.62 0.47
Right 0.51 0.39
Reliability Most-unreliable 0.55 0.42
Unreliable 0.61 0.47
Reliable 0.64 0.49
Most-reliable 0.57 0.42
Deep interaction rates: We next consider the correlation between the deep interactions ratio for combined (DIRC) and public (DIRP) posts. Table 4 summarizes these results.
First, when comparing the correlations for deep interaction ratios across the bias classes, we observe that the Left and Center bias classes show higher correlations across both metrics, while the Right class exhibits weaker relationships. This pattern suggests that deep interactions, or more in-depth engagements with content, are more consistent between public and private contexts for left-leaning and centrist articles. Second, in terms of reliability, the correlations are relatively similar across classes, with the Reliable class showing the strongest relationships and the Most-unreliable class having the weakest. These findings indicate that the consistency of deep interactions across public and private contexts is less affected by articles’ perceived reliability than their political bias.
7 Discussion
Our research provides a comprehensive exploration of news interactions on Facebook, with a specific focus on discerning the nuances between public and combined interactions across diverse news classes. We developed a robust comparison methodology, crucial for future studies in this area. Our findings, presented in several key facets, provide valuable insights along several dimensions (discussed next).
Reliability and Bias: Several discoveries emerged when examining the reliability of news articles. Surprisingly, articles classified as Reliable garnered lower levels of engagement compared to those classified as Most-reliable or even various Unreliable categories. Furthermore, our investigation into the impact of biases unveiled that biased news, regardless of whether it leans Left or Right, consistently received higher levels of user interaction than more neutral (Center) content.
Public vs. Private: One of the central aspects of our study concerned the comparisons of the interaction patterns within the public vs. private spheres. What became evident was that articles classified as Most-reliable or Most-unreliable, representing the extremities of the reliability spectrum, attracted significantly higher proportions of engagement within private spaces compared to their public counterparts. Intriguingly, no statistically significant differences were observed between these two extreme groups here. This observation holds substantial implications for content moderators, policymakers, and researchers, prompting us to ponder whether an exclusive focus on the public facet of Facebook provides a comprehensive understanding of user interactions, particularly concerning news content classified as Most-unreliable.
Interaction Depth: Our research extended beyond interaction frequency to explore the depth of interactions, revealing noteworthy variations between the public and private spheres. Notably, statistically significant differences in interaction depth were observed, reinforcing the importance of considering both spheres in research endeavors. For example, the extremities of the reliability spectrum (Most-reliable and Most-unreliable) exhibited a tendency to elicit deeper interactions from users within the public sphere. However, this narrative shifted when examining combined interactions, indicating that only the Most-unreliable class triggered statistically significant deeper engagement from Facebook users. This underscores the intricate interplay between reliability, bias, and interaction depth, urging researchers to explore these dynamics beyond the confines of public engagement.
Ethical Considerations: It is essential to acknowledge the ethical dimension of our research. While our findings offer invaluable insights, they also hold the potential for misuse. To mitigate this risk, we undertook several precautions, such as selecting a diverse set of articles and news sources and maintaining objectivity in our analysis. We adhered to relevant privacy regulations, ensuring that no personally identifiable information was collected or disclosed. Transparency is of utmost importance to us, and we will provide all codes used in this paper, along with the list of news articles, in a camera-ready version. Although we cannot publish raw engagement data due to CrowdTangle terms of service, we will furnish attachment status information for URLs and posts through a lookup table with hashed keys. This approach enables responsible utilization of our research findings while minimizing the likelihood of negative consequences.
Beneficiaries: Our findings hold value for a diverse range of beneficiaries. Researchers, journalists, and policymakers can gain insights into news engagement patterns, aiding in informed decision-making. Social media content moderators can benefit by understanding where to focus their efforts. Moreover, our insights have relevance for platforms and media organizations, guiding them in shaping content strategies for more effective audience engagement.
In conclusion, our research sheds light on complex engagement dynamics on social media and emphasizes the responsible use of these findings for informed decision-making.
8 Related Works
This research most closely relates to studies exploring user engagement dynamics with various social media content. Our main contribution lies in the novel analysis of interaction disparities between public and private news article sharing on Facebook, particularly concerning articles with varying bias and reliability. Examining interactions depth provides another level of depth to our analysis and findings.
There exist studies that partially touch upon some dimensions highlighted in this paper. However, as discussed in the introduction, the vast majority of these works, particularly pertaining to Facebook, predominantly concentrate on the platform’s public sphere, neglecting to provide a comparative analysis with the private or combined sphere. To contextualize our study within the context of the existing literature and to highlight the novelty of our contribution, we next describe the most closely related works.
One line of this prior research investigates user engagement with different classes of news, particularly regarding bias and reliability. For example, Edelson et al. [7], while focusing on the public sphere, conducted a large-scale study on user engagement with 7.5 million posts from 2551 publishers (Facebook pages) across both bias and reliability dimensions. Their results show that individual posts from non-misinformation news outlets tend to attract lower median engagement than misinformation. Weld et al. [34] carried out a similar investigation on Reddit rather than on Facebook. When considering the reliability parameter, Vosoughi et al. [35] analyzed a dataset of approximately 126,000 stories tweeted by over 3 million people more than 4.5 million times, finding that false news stories reached more people than true news and diffused significantly farther, faster, deeper, and more broadly than the truth across various categories of information. Notably, during the critical period of the COVID-19 pandemic and by doing a cross-national study, Altay et al. [36] delineated the surge in online news consumption, with credible news outlets witnessing a significant boost. Samory et al. [37] characterized the social media news sphere through user co-sharing practices by focusing on 639 news sources, both credible and questionable, and divided them into 4 clusters, and characterizing them according to the audience that shares their articles on Twitter and how the stylometric features used by each cluster is successful in getting users engagement. Lamot et al. [38] examined how news headlines are remediated on Facebook and how this affects user engagement. Finally, Boukes et al. [39] compared user-content interactivity and audience diversity across news and satire, finding differences in online engagement between satire, regular news, and partisan news. Considering the bias dimension, on the other hand, Wischnewski et al. [40] discovered that users are more inclined to share hyperpartisan news articles that coincide with their own political views. The bias towards the right party in sharing and engagement with news is another thesis that has been studied by previous works on both Twitter [41] and Facebook [42].
Other dimensions of user engagement have also been considered. For example, Aldous et al. [8] investigated the effects of topic and emotional factors on posting across five social media platforms (Facebook, Instagram, Twitter, YouTube, and Reddit). They found that user engagement is strongly influenced by the content’s topic, with certain topics being more engaging on a particular platform. By controlling the effect of the author (considering articles by the same author) over a one-year period, Maier et al. [43], among other things, found that highly emotional topics get a higher level of Facebook engagement. A large body of research has also focused on studying the role of sad and anger in social media news sharing and engagement (e.g., [44, 45]).
Finally, we note that profiling of the relation between user engagement dynamics and different content itself has paved the way for new applications. Examples include techniques and models that rely (partly) on user engagement dynamics (as features) for the detection of different content types. Such applications may benefit social media content moderators [46, 47]. Others have profiled the temporal dynamics of user engagement with different classes of news [12, 48].
In summary, while previous research has partly explored some dimensions of user engagement with news articles, our study is the first to directly compare the dynamics of public and private sharing of news articles on Facebook, specifically examining the interactions and depth of engagement with articles of different levels of bias and reliability.
9 Conclusions
In this paper, we presented the first investigation of the differences in interaction patterns between public and private sharing of news articles on Facebook, focusing on articles with varying bias and reliability, as well as the depth of interactions. By analyzing a dataset of over 19K news articles (with more than 225M interactions on Facebook), employing the CrowdTangle browser addon for data collection, and carefully processing this data, we provided a comprehensive analysis of user interaction with news content on Facebook.
Our findings reveal significant differences in interaction patterns between public and private sharing across different classes of news. Notably, users tend to engage more deeply in private discussions compared to public ones, irrespective of the bias or reliability class of the news. We also found that Facebook users engage more deeply with the least reliable news and the right-biased news, both classes exhibiting higher deep interaction ratio levels. While this tendency may suggest that the content from these classes may be more polarizing or controversial, leading to deeper discussions and debates, further research (e.g., by interviewing users) is needed to determine the underlying factors driving this difference in engagement.
Our research findings offer valuable insights with significant implications for various stakeholders (e.g., researchers, journalists, policymakers, and social media content moderators) and deepen our understanding of how public and private spheres influence the dissemination and engagement with news articles on social media, enabling more informed decision-making. For example, our identification of content classes that generate deeper user engagement can provide actionable guidance to Facebook moderators, enabling them to focus their efforts effectively. Furthermore, the disparities we uncovered between public and private interaction patterns can guide future research that considers both spheres, promoting a more comprehensive approach. To facilitate such analysis, we present a robust methodology developed through active collaboration with CrowdTangle, which ensures a fair comparison between public and private spheres. This methodology equips researchers and content moderators with a valuable tool to navigate the evolving landscape of online news engagement.
In conclusion, by examining the differences in public and private sharing of news articles, this study offers valuable insights into the user engagement patterns associated with contents belonging to different bias and reliability classes, as well as highlights the importance of understanding the dynamics of news sharing and interaction both in the public and private spheres of social media platforms such as Facebook. While challenges remain, our methodology and discoveries offer valuable data-driven guidance toward addressing both spheres.
10 Statements and Declarations
10.1 Conflict of interest
The authors have no competing interests.
10.2 Author contributions
Both authors have made substantial contributions to the conceptualization, design, data collection, analysis, and writing of the research.
10.3 Acknowledgment
The authors thank the CrowdTangle team for their active feedback and help during this research.
10.4 Data and code availability
All the codes associated with this research will be made available with the camera-ready version. At the same time, to facilitate replicating the results and conducting similar research, the list of all news articles considered in this study and also the attachment status of all related posts will be published.

### v3

@@
-Understanding Differences in News Article Interaction Patterns on Facebook: Public vs. Private Sharing with Varying Bias and Reliability
-Abstract
-The rapid growth of news dissemination and user engagement on social media has raised concerns about the influence and societal impact of biased and unreliable information.
-As a response to these concerns, a substantial body of research has been dedicated to understanding how users interact with different news.
-However, this research has primarily analyzed publicly shared posts.
-With a significant portion of engagement taking place within Facebook’s private sphere, it is therefore important to also consider the private posts.
-In this paper, we present the first comprehensive comparison of the interaction patterns and depth of engagement between public and private posts of different types of news content shared on Facebook.
-To compare these patterns, we gathered and analyzed two complementary datasets: the first includes interaction data for all Facebook posts (private + public) referencing a manually labeled collection of over 19K news articles, while the second contains only interaction data for public posts tracked by CrowdTangle.
-As part of our methodology, we introduce several carefully designed data processing steps that address some critical aspects missed by prior works but that (through our iterative discussions and feedback with the CrowdTangle team) emerged as important to ensure fairness for this type of study.
-Our findings highlight significant disparities in interaction patterns across various news classes and spheres.
-For example, our statistical analysis demonstrates that users engage significantly more deeply with news in the private sphere compared to the public one, underscoring the pivotal role of considering both the public and private spheres of Facebook in future research.
-Beyond its scholarly impact, the findings of this study can benefit a wide range of stakeholders, including Facebook content moderators, regulators, and policymakers, contributing to a healthier online discourse.
-keywords:
-News sharing, Facebook, User engagement, Bias, Reliability, Public vs private
-1 Introduction
-Social media has become a dominant medium for news consumption and sharing, with Facebook being the most popular and widely used platform for these purposes [1, 2].
-For example, according to a 2021 Pew Research Center survey [2], 66% of all Americans regularly use Facebook, and almost half of them (31% of the US population) regularly get their news via the platform.
-Other popular platforms that users regularly use for news consumption include YouTube (22% of the US population), Twitter (13%), and Instagram (11%).
-This significant reach highlights the potential impact that news consumption and sharing on Facebook (and social media in general) can have on the opinions and thoughts of both individuals and the society [3].
+Public versus Less-Public News Engagement on Facebook: Patterns Across Bias and Reliability
+Abstract.
+The rapid growth of social media as a news platform has raised significant concerns about the influence and societal impact of biased and unreliable news on these platforms.
+While much research has explored user engagement with news on platforms like Facebook, most studies have focused on publicly shared posts.
+This focus leaves an important question unanswered: how representative is the public sphere of Facebook’s entire ecosystem?
+Specifically, how much of the interactions occur in less-public spaces, and do public engagement patterns for different news classes (e.g., reliable vs.
+unreliable) generalize to the broader Facebook ecosystem?
+This paper presents the first comprehensive comparison of interaction patterns between Facebook’s more public sphere (referred to as public in paper) and the less public sphere (referred to as private).
+For the analysis, we first collect two complementary datasets: (1) aggregated interaction data for all Facebook posts (public + private) for 19,050 manually labeled news articles (225.3M user interactions), and (2) a subset containing only interactions with public posts (70.4M interactions).
+Then, through discussions and iterative feedback from the CrowdTangle team, we develop a robust method for fair comparison between these datasets.
+Our analysis reveals that only 31% of news interactions occur in the public sphere, with significant variations across news classes.
+Engagement patterns in less-public spaces often differ, with users, for example, engaging more deeply in private contexts.
+These findings highlight the need to examine both public and less-public engagement to fully understand news dissemination on Facebook.
+The observed differences hold important implications on content moderation, platform governance, and policymaking, contributing to healthier online discourse.
+journalyear: 2025copyright: ccconference: Proceedings of the 17th ACM Web Science Conference 2025; May 20–24, 2025; New Brunswick, NJ, USAbooktitle: Proceedings of the 17th ACM Web Science Conference 2025 (Websci ’25), May 20–24, 2025, New Brunswick, NJ, USAdoi: 10.1145/3717867.3717912isbn: 979-8-4007-1483-2/25/05ccs: Information systems World Wide Web
+1. Introduction
+Social media has emerged as an increasingly important news source, with Facebook maintaining a prominent position among social platforms  (Newman et al., 2023; Christopher Aubin, 2024).
+For example, according to a 2024 Pew Research Center report (Christopher Aubin, 2024), 33% of U.S.
+adults regularly get news on Facebook, similar to YouTube (32%), and notably ahead of Instagram (20%) and Twitter (12%).
+This extensive reach highlights the potential impact of news sharing on Facebook in shaping both individual opinions and broader societal discourse (Thompson et al., 2020).
 Unfortunately, not all news are reliable.
-With an increasing amount of misinformation being circulated on Facebook, it is crucial to understand how users engage with news of varying reliability [4].
-Improved insights and understanding into these dynamics are expected to benefit media researchers, journalists, content moderators, and policymakers, which choices based on these insights, in turn, are expected to influence public opinion and the behavior of regular users [5, 6].
-Previous research has extensively examined public user engagement dynamics with different news content and the factors affecting this engagement  [7, 8, 9, 10, 11, 12].
-However, with a growing privacy inclination of Facebook users and a rise in user engagement in private spaces of Facebook [13, 14], it is becoming increasingly important to understand these dynamics also with regards to private sharing.
-Yet, the current literature predominantly focuses on public posts, leaving gaps in our understanding of engagement differences between these realms.
-Most importantly, prior research has not studied the variations in public vs.
-non-public interaction with regards to (1) reliable and unreliable news content, (2) news with different political biases, or (3) the depth of the interactions associated with each of these types of posts and content types.
-These are the primary gaps addressed in this paper.
-In this paper, we examine the differences in public and private sharing of news articles on Facebook, with an emphasis on users’ interactions with these articles.
-Specifically, we investigate how news articles written with different bias and reliability are shared and engaged with, as well as the depth of these interactions.
-In this context, bias refers to a tendency for news articles to exhibit partiality or favoritism towards particular groups or ideas (e.g., left or right on the political spectrum), while reliability refers to the accuracy and credibility of the information presented [15] (e.g., fake or true).
-While most previous research focuses on either bias or reliability, we consider both dimensions, as they have been found related, but yet each has its own ability to influence news article sharing behaviors and affect the quality and diversity of information that users encounter on social media [16, 17].
-On the other hand, by considering whether an interaction is deep or shallow, we can identify whether there are areas (private or public) where users engage more deeply with specific content types (e.g., fake news).
-This is important since the depth of user engagement can significantly influence the success of an article.
-For example, an article may have greater success if people share and comment on the posting of the article than if they only press Facebook’s “like” button [1].
-Research questions and methodology: With the above goals in mind, we designed our study to address the following research questions:
+With an increasing amount of misinformation being circulated on Facebook, it is crucial to understand how users engage with news of varying reliability (Geeng et al., 2020).
+Understanding of these dynamics are expected to benefit media researchers, journalists, content moderators, and policymakers, whose choices based on these insights, in turn, are expected to influence public opinion and the behavior of regular users (Nah and Yamamoto, 2018; Jones et al., 2017).
+Previous research has extensively examined public user engagement dynamics with different news content and the factors affecting this engagement  (Edelson et al., 2021; Aldous et al., 2019; Brena et al., 2019; Marwick, 2018; Vitak et al., 2011).
+However, with a growing privacy inclination of Facebook users and a rise in user engagement in private spaces of Facebook (Dey et al., 2012; Lottridge and Bentley, 2018), it is becoming increasingly important to understand these dynamics also with regards to private sharing.
+Yet, the current literature is centered around more public posts, leaving gaps in our understanding of engagement differences between these realms.
+Most importantly, prior research has not studied the dynamics of public vs.
+private interaction.
+To address this gap, in this paper we present the first comprehensive comparison of the news article sharing and user interaction patterns seen on Facebook’s more public sphere (referred to as public in paper) versus in the less public sphere (for simplicity referred to as private).
+Specifically, we investigate differences in how news articles written with different bias and reliability are shared and engaged with, as well as the depth of these interactions.
+In this context, bias refers to a tendency for news articles to exhibit partiality or favoritism towards particular groups or ideas (e.g., left or right on the political spectrum), while reliability refers to the accuracy and credibility of the information presented (e.g., fake or true).
+While most previous research on news engagement has primarily focused on either bias or reliability (and on public engagement), we consider both dimensions, as they have been found related, but yet each has its own ability to influence news article sharing behaviors and affect the quality and diversity of information that users encounter on social media (Guess et al., 2018; Pennycook et al., 2018).
+Furthermore, by analyzing whether interactions are deep or shallow, we can determine if users engage more deeply with certain content types (e.g., fake news) in private or public spaces.
+This distinction is crucial, as deeper engagement—such as sharing or commenting—can significantly enhance an article’s impact compared to shallow actions like pressing Facebook’s “like” button (Newman et al., 2023).
+Research questions and methodology: Guided by the goal of better understanding how representative the public sphere of Facebook’s ecosystem (typically studied in prior works) is of the private sphere and the ecosystem as a whole, we designed our study to address the following research questions:
   1. RQ1
@@
 We took several steps to address these previously unaddressed questions in as controlled matter as possible.
-First, in contrast to most prior works that use publisher-level labeling of news (e.g., Adelson et al.
-[7]; Horne et al.
-[18]), we employ article-level labeling of bias and reliability.
-By doing so, we capture that not all articles by a publisher have the same bias or reliability, recognizing that the publisher (source) is only one factor [19] in the bias/reliability.
+First, in contrast to most prior works that use publisher-level labeling (e.g., Adelson et al.
+(Edelson et al., 2021); Horne et al.
+(Horne et al., 2019)), we employ article-level labeling of bias and reliability.
+By doing so, we capture that not all articles by a publisher have the same bias or reliability, recognizing that the publisher (source) is only one factor (Pidikiti et al., 2020) in the bias/reliability.
 In particular, we consider the bias and reliability of articles ($N=19,050$), meticulously selected from a substantial pool of publishers ($N=1,121$) and 30K+ manually labeled news articles.
-Here, the original labeling was provided by Ad Fontes Media [20], which rates both news sources and individual articles based on their level of bias and reliability.
+Here, the original labeling was provided by Ad Fontes Media (Ad Fontes Media, 2024), which provides bias and reliability of news articles.
 Second, through active back-and-forth discussions and feedback with the Crowdtangle team, we developed and implemented a methodology using their Chrome addon that allows us to obtain simultaneous interaction statistics for a representative sample of publicly shared Facebook posts linking these articles, as well as across all Facebook posts (including both public and private posts) linking these articles.
@@
 Our results also show that reliable news content has significantly lower private interaction shares compared to the highly-reliable or even unreliable content.
-Example beneficiaries: Our methodology and findings contribute to a better understanding of the dynamics of news sharing and interaction on Facebook and provide valuable insights for multiple stakeholders.
-For example, considering the significant impact that social media interactions have on content and platform popularity [21], our findings on how bias and reliability influence user engagement in public vs.
-private spheres can aid media researchers and journalists.
-These insights can inform more effective news content creation and distribution strategies for enhanced audience engagement.
-For content moderators of platforms such as Facebook, the analysis offers data-driven guidance on where moderation efforts may be most needed to curb the spread of problematic content.
-For policymakers and regulators, our findings illustrate how user privacy settings impact news content engagement types and interaction depth.
-Notably, highly unreliable news receives more private discussion engagement share compared to the broader unreliable category.
+Example beneficiaries: Our methodology and findings contribute to a better understanding of news sharing and interaction dynamics on Facebook, offering valuable insights for various stakeholders.
+Media researchers and journalists can use our analysis of how bias and reliability shape engagement in public and private spheres for improved content creation and distribution strategies.
+For content moderators, the analysis offers data-driven guidance on prioritizing efforts to curb the spread of problematic content.
+For policymakers, we highlight how privacy settings influence engagement patterns, with highly unreliable news, for example, garnering more engagement in the private sphere.
 These results underscore how platform design choices can influence selective exposure and engagement, potentially exacerbating issues like ideological polarization and misinformation spread.
 Roadmap: Sect.
-2 explains our multi-faceted research design, including news article selection, labeling, interactions data collection, and the processing steps taken to ensure fair comparisons between public and private engagement.
+2 explains our research design, including news article selection, labeling, data collection, and the processing steps taken to ensure fair comparisons.
 The resulting dataset is summarized in Sect.
@@
 5, we turn our attention to differences in the depth of interactions.
-Sect.
-6 concludes our analysis with a correlation-based analysis between public and private interactions.
-Discussion of the key findings is provided in Sect.
-7.
+Topic and top-publishers analysis are provided in Sects.
+6 and 7.
 Finally, related works (Sect.
@@
 9) are presented.
-2 Research Design, Data Collection and Limitations
+2. Research Design, Data Collection and Limitations
 At a high level, our research design has three parts.
-  • 1.
-    Article selection and labeling: We first obtain the URLs to a substantial number of news articles, which have been classified by Ad Fontes Media [20] with regards to both their individual bias and reliability.
-We then decide category thresholds and label each article into one of three bias classes and one of four reliability classes.
-  • 2.
-    Collection of interactive data: After careful preprocessing of the URLs, we use the CrowdTangle browser extension to obtain interaction data for two sets of posts linked to these URLs (1) public and (2) combined (public + private), where the first set is a subset of the second.
+Part 1.
+Article selection and labeling: We first obtained bias and reliability scores for the articles evaluated by Ad Fontes Media (Ad Fontes Media, 2024).
+Subsequently, each article was categorized into one of three bias classes and one of four reliability classes based on these scores.
+Part 2.
+Collection of interactive data: After careful preprocessing of the URLs, we used the CrowdTangle browser extension to obtain interaction data for two sets of posts linked to these URLs (1) public and (2) combined (public + private).
 To address some limitations of the API and to ensure fair comparison of the sets, we leave a four-month gap between the latest labeled article (Nov.
 1, 2022) and the primary data collection (Mar.
-2023), as well as apply some additional post-processing (e.g., examining the actual posts), filtering (e.g., based on thresholds determined via communication with the CrowdTangle team), and collect some complementing data directly from CrowdTangle (e.g., to address limitations of the extension API).
-  • 3.
+2023), as well as apply some additional post-processing (e.g., examining the actual posts), filtering (based on thresholds determined via discussions with the CrowdTangle team), and collect some complementing data directly from CrowdTangle (to address limitations of the extension API).
+Part 3.
 Data analysis: Finally, we use the final dataset (capturing the two sets) to compare the properties of the public vs.
-combined (private + public) sets and discover statistically significant differences between the interaction patterns of public and private posts.
-Here, it should also be noted that the combined set is dominated by private posts.
-We next provide detailed descriptions of the initial two steps, before presenting our analysis results (step 3) in the subsequent sections.
-2.1 News Article Selection and Labeling
+combined (private + public) sets and identify significant statistical differences between the interaction patterns of public and private posts.
+We next provide details of the initial two steps, before presenting our analysis results (Part 3) in the subsequent sections.
+2.1. News Article Selection and Labeling
 Selection of articles: There are several independent initiatives that assess the bias and/or reliability of individual news articles and/or news sources.
-Examples of such evaluation efforts include Media Bias Fact Check [22], Ad Fontes Media, AllSides [23], and NewsGuard [24].
-Among these, we opted to use data from Ad Fontes Media for the following main reasons: (1) it has been widely used in previous research [25, 26, 27, 12], (2) they evaluate individual news articles (not only the news publishers), (3) each assessed article receives a score for both bias and reliability, (4) they offer a transparent evaluation methodology, which is published and explained in a white paper [28], and finally (5) the dataset comprises a large set of news contents (including news articles, podcasts and TV shows) randomly selected from a diverse set of publishers.
+Examples of such evaluation efforts include Media Bias Fact Check (Media Bias Fact Check, 2024), Ad Fontes Media, AllSides (AllSides, 2023), and NewsGuard (NewsGuard, 2024).
+Among these, we opted to use data from Ad Fontes Media for the following main reasons: (1) it has been widely used in previous research (Haq et al., 2022; Huszár et al., 2022; Knutson et al., 2024), (2) they evaluate individual news articles (not only the news publishers), (3) each assessed article receives a score for both bias and reliability, (4) they offer a transparent evaluation methodology, which is published and explained in a white paper (Otero, 2021), and finally (5) the dataset comprises a large set of news contents randomly selected from a diverse set of publishers.
 We obtained all evaluated news from Ad Fontes Media as of Nov.
 1, 2022.
-After filtering the podcasts and TV shows, this set included 31,446 news article URLs from 1,121 publishers.
-We also filtered a few URLs that (1) primarily served as event reporting pages (e.g., https://www.nola.com/news/hurricane) and (2) articles for which Ad Fontes Media had changed their rating (identified through complementary documents published by Ad Fontes Media).
-After excluding these articles, the dataset comprised 31,408 articles.
+This included 31,446 news article URLs from 1,121 publishers.
+After filtering out URLs serving as event reporting pages (e.g., https://www.nola.com/news/hurricane) and articles with updated ratings (per Ad Fontes Media’s documentation), the final dataset included 31,408 articles.
 Labeling of articles: Each article in the Ad Fontes Media dataset is assessed for both bias and reliability by a minimum of three human analysts, representing a mix of right, left, and center self-reported political perspectives.
@@
 As for reliability, scores vary from 0 to 64, with 64 representing the most-reliable news.
-Figure 1: CDF of bias scores
-Figure 2: CDF of reliability scores
 For our analysis, we categorized the bias and reliability scores into distinct classes.
@@
 True), we opted for a more granular approach by defining three classes of bias (Left, Right, and Center) and four classes of reliability (Most-unreliable, Unreliable, Reliable, and Most-reliable).
-Figures 1 and 2 show the Cumulative Distribution Function (CDF) of the bias and reliability scores of the articles included in our analysis (after additional filtering steps, explained later in this section) together with the threshold values and class labels used to define each class.
+Figures 1(a) and 1(b) show the Cumulative Distribution Function (CDF) of the bias and reliability scores of the articles included in our analysis (after additional filtering steps, explained later in this section) together with the threshold values and class labels used to define each class.
 With our labeling, anything with a bias score below -6 is called Left-biased, and anything above 6 is called Right-biased, with the range (-6,6) capturing the Center class.
-We note that Ad Fontes Media also refers to this range as “Middle or Balanced Bias” [20].
-As observed, this class accounts for most samples in our dataset, representing 58% (11,094 out of 19,050 articles in the final dataset).
+We note that Ad Fontes Media also refers to this range as “Middle or Balanced Bias” (Ad Fontes Media, 2024).
+As observed, this class represents most samples in our dataset, representing 58% (11,094 out of 19,050 articles in the final dataset).
 With this split, the two biased classes have similar, close to 20%, shares.
 For the reliability classes, we assigned the Most-unreliable class to cover the lowest range (0, 32).
-According to Ad Fontes Media’s terminology [20], these values typically correspond to articles “which contain inaccurate and misleading info, selective stories or have opinion and wide variation in their reliability”.
+According to Ad Fontes Media’s terminology (Ad Fontes Media, 2024), these values typically correspond to articles “which contain inaccurate and misleading info, selective stories or have opinion and wide variation in their reliability”.
 Similarly, the range (48-64), chosen for the Most-reliable class, encompasses articles characterized by “thorough and original fact reporting” due to Ad Fontes terminology.
 Finally, the threshold of 40, between Reliable and Unreliable, matches the mid-point between these two classes as well as where Ad Fontes make their split between “wide variation in reliability” and “mix of fact reporting and analysis”.
-2.2 Collection of Interaction Data
-Preprocessing of URLs: Before collecting the interaction data associated with each article (URL), we (1) replaced the link shorteners with their corresponding URLs and (2) converted all article URLs to their canonical forms.
-During the conversion process, we primarily removed all unnecessary URL parameters (those sent after the ”?” symbol) except for the ones essential to the canonical form (e.g., “id” parameters are sometimes part of the canonical form).
-As another step in identifying the canonical forms, we replaced all the web archive links in the Ad Fontes dataset with the original links.
-Primary data collection for the two sets of posts: We used the CrowdTangle browser extension [29] to collect the interaction data for the Facebook posts sharing each news article of interest.
-In addition to interaction statistics for the public posts that CrowdTangle (owned by Facebook) indexes and tracks over time (their database, available for and used by researchers, contains longitudinal interaction data for indexed posts), this extension provides access to the life-time interaction statistics, obtained thus far, across all posts (private + public).
-In this paper, we use this extension to collect statistics about all interactions associated with these two sets of posts: combined and public, where the combined captures all interactions of all posts (public + private) and the public is a subset of the combined (only capturing public posts).
+2.2. Collection of Interaction Data
+Preprocessing of URLs: Before collecting the interaction data associated with each article (URL), we (1) expanded link shorteners and (2) converted URLs to canonical forms.
+During the conversion process, we primarily removed unnecessary URL parameters after “?” except those essential to canonical forms (e.g., “id” parameters are sometimes part of the canonical form).
+We also replaced web archive links in the Ad Fontes Media’s dataset with original links.
+Primary data collection for the two sets of posts: Prior to its shutdown in Aug.
+2024, CrowdTangle (owned by Facebook) provided access to interaction statistics for public posts, which it indexed in a database widely used by researchers.
+Additionally, their Chrome extension offered access to the life-time interaction statistics available at the time, encompassing all posts (public + private).
+In this study, we utilized the extension to collect statistics for two datasets: combined and public, where the combined dataset includes all interactions from all posts (public + private), and the public dataset represents a subset of the combined dataset, containing only interactions with publicly accessible posts.
+(a) Bias scores
+(b) Reliability scores
+Figure 1. CDFs of scores for (a) bias and (b) reliability.
 To collect the data for each URL, we used the “Download” option within the extension’s interface when browsing each article.
-This allowed us to easily download a CSV file containing the interactions data for both sets.
+The downloaded CSV file contained the interaction data for both sets.
 As expected from a privacy standpoint, only the aggregated statistics are provided for the combined set.
-In contrast, every post and their individual interaction data are provided for the public set.
+In contrast, every post and its individual interaction data are provided for the public set.
 At a high level, the combined category aggregates the interactions (e.g., likes, shares, comments, etc.) of all Facebook posts referencing the article’s URL, regardless of the post’s privacy settings.
@@
 In contrast, the public category, includes all interactions with the public posts tracked by CrowdTangle, which, due to heavy-tailed characteristics and significant coverage, captures most interactions with public Facebook posts.
-As supporting examples, all posts of all US-based public groups with 2K+ members are indexed, and in 2021 over 99% of all Facebook pages with at least 25K likes were indexed [30].
-While some public posts are not indexed, we note that their indexing does not consider any bias/reliability criteria, is expected to capture the majority of public posts, and the majority of the posts in the combined set are private.
+As supporting examples, all posts of all US-based public groups with 2K+ members are indexed, and in 2021 over 99% of all Facebook pages with at least 25K likes were indexed (Tess, 2021).
 Timeline and limitations: We collected the interaction data in Mar.
 2023, ensuring at least a four-month gap between the publication date of the newest article in our dataset and our interaction data collection.
-This length of interval was chosen to ensure that we capture most of the posts and interactions for the URLs in our dataset.
-As mentioned in [31] and confirmed via our communication with CrowdTangle, the information from the CrowdTangle addon originates from two different sources: (1) public interaction data is obtained from the CrowdTangle database, while (2) combined interaction data is derived from Facebook’s Graph API.
-Consequently, variations in data capturing times might cause some URLs to exhibit higher public than combined interactions [31].
-Furthermore, as discussed in [31], Facebook’s aggregation data restarts the interaction count whenever websites update their connection schemes (e.g., switching from HTTP to HTTPS) after publishing a URL, while CrowdTangle continues accumulating data.
+This interval ensured capturing most posts and interactions for the URLs.
+This four-month window is notably conservative, as research has shown that most interactions occur within days of publication (Thorgren et al., 2024; Pfeffer et al., 2023).
+Although this conservative window may have overlooked interactions with content that was later removed or deleted, we are unable to assess the extent of this effect—if it exists—due to CrowdTangle’s shutdown in Aug.
+2024.
+As mentioned in (Garmur, 2022), the information from the CrowdTangle addon originated from two different sources: (1) public interaction data is obtained from the CrowdTangle database, while (2) combined interaction data is derived from Facebook’s Graph API.
+Consequently, variations in data capturing times might cause some URLs to exhibit higher public than combined interactions.
+Furthermore, as confirmed by the CrowdTangle team, Facebook’s aggregation data restarts the interaction count whenever websites update their connection schemes (e.g., switching from HTTP to HTTPS) after publishing a URL, while CrowdTangle continued accumulating data.
 Although the discrepancy is not substantial (9% in our initial dataset), we aimed to mitigate the effect of these differences where possible (described next).
-Further filtering and data processing for fairer comparisons: After consulting with the CrowdTangle team regarding all numerical discrepancies that we observed, we identified two additional reasons for the discrepancies.
-First, as stated in [32], for privacy reasons, for the combined data (that include private posts) “the values are intentionally not precise, but you can be confident they accurately reflect user engagement with a URL.” Discussing the identified discrepancy for such cases, the CrowdTangle team confirmed that combined interaction data with counts below 100 is not reliable; however, that data with counts above 100 provides a reliable estimate of public + private interactions.
-For this reason, we removed all URLs in our dataset with combined interactions below 100.
-Although this reduced our dataset to 19,505 articles, it is important to note that this group of URLs accounted for less than 4% of public interactions in our dataset.
-Furthermore, we note that the public interaction data for this group still remains reliable.
-Second, the CrowdTangle team informed us that the Facebook Graph API only considers posts with the URL attached to the post (i.e., those displaying a preview of the URL) when counting interaction with the URL, while CrowdTangle’s algorithm includes any post containing a linked URL, regardless of attachment status.
-Taking this into account, in our final step toward enhancing interaction data quality, we chose not to rely on the aggregated data from the CrowdTangle extension (which sums the interactions of all posts mentioned in the extension).
-Instead, we examined each post to determine whether the related URL was attached or not, and if so, we included them in the sum of public interactions for that URL.
-[^1] The attachment status of all the posts and URLs will be published with the camera-ready version for the interested researchers.
-This resulted in a more accurate comparison between combined and public interactions.
-After these processing steps, we had only 455 URLs for which public interactions exceeded combined interactions, which were excluded from our final dataset.
-The primary causes of these discrepancies were beyond the scope of our research to address (e.g., changes in URL, publisher protocol schemes, or different data capturing times as mentioned above).
-While some of the pre- and post-processing steps described above require significant effort, they help ensure an accurate and fair comparison between the public and combined sets in such a way that we can provide conclusive insights into the relative sharing patterns of public vs.
-private posts.
-Enhancing the data with additional posts: Another limitation of the CrowdTangle extension is that it only allows us to retrieve data for up to 500 posts per URL.
-To address this limitation, for the URLs with more than 500 posts in the public interaction category, we obtained additional data directly from CrowdTangle.
-However, these URLs constitute only 42 instances in our dataset and will have no significant impact on the reproducibility of our statistical analysis.
-3 Dataset
-High-level summary: After applying the aforementioned filtering steps, we have in total 19,050 articles from 1,121 news outlets remaining in our dataset.
-These articles have been shared in 253,350 posts, which combined are responsible for 225,282,745 interactions (out of which 70,428,316 are public).
-Table 1 summarizes the dataset.
-Bias data: Table 2 shows the number of articles and the total combined and public interactions for each bias and reliability class.
-With the selected bias thresholds, our dataset includes 4,516 Left-biased articles and 3,440 Right-biased articles, with the remaining 11,094 articles falling in the Center category.
-These articles are in turn responsible for 79,038,246 (Left), 56,160,101 (Right), and 90,084,398 (Center) interactions.
-While the Center category includes 58% of the articles, it is interesting to note that it is responsible for a significantly smaller fraction of the total interactions (40%).
-Instead, the Left-biased and Right-biased articles see relatively higher interaction rates (studied in the next section).
-For example, the Left-biased articles account for only 24% of the articles but 35% of the interactions, and the Right-biased articles are responsible for only 18% of the articles but 25% of the interactions.
-Table 1: Summary of the final dataset
-Type Articles Publishers Posts Combined interactions Public interactions
-Number 19,050 1,121 253,350 225,282,745 70,428,316
-Table 2: Statistics based on Bias and Reliability
+Table 1. Statistics based on Bias and Reliability.
 Class Articles no Combined interactions Public interactions
@@
 Most-reliable 2,427 43,674,206 12,382,378
+Total 19,050 225,282,745 70,428,316
+Figure 2. Interactions per article
+Further filtering and data processing for fairer comparisons: After consulting with the CrowdTangle team regarding all numerical discrepancies that we observed, we identified two additional reasons for the discrepancies.
+First, as stated in (Meta, 2023), for privacy reasons, for the combined data (that include private posts) “the values are intentionally not precise, but you can be confident they accurately reflect user engagement with a URL.” Discussing the identified discrepancy for such cases, the CrowdTangle team confirmed that combined interaction data with counts below 100 is not reliable; however, that data with counts above 100 provides a reliable estimate of public + private interactions.
+For this reason, we removed all URLs in our dataset with combined interactions below 100.
+Although this reduced our dataset to 19,505 articles, it is important to note that this group of URLs accounted for less than 4% of public interactions in our dataset.
+Furthermore, we note that the public interaction data for this group still remains reliable.
+Second, the CrowdTangle team informed us that the Facebook Graph API only considers posts with the URL attached to the post (i.e., those displaying a preview of the URL) when counting interaction with the URL, while CrowdTangle’s algorithm includes any post containing a linked URL, regardless of attachment status.
+Taking this into account, in our final step toward enhancing interaction data quality, we chose not to rely on the aggregated data from the CrowdTangle extension (which sums up the interactions of all posts mentioned in the extension).
+Instead, we examined each post to determine whether the related URL was attached or not, and if so, we included them in the sum of public interactions for that URL.
+This resulted in a more accurate comparison between combined and public interactions.
+After these processing steps, we had only 455 URLs for which public interactions exceeded combined interactions, which were excluded from our final dataset.
+The primary causes of these discrepancies were beyond the scope of our research to address (e.g., changes in URL, publisher protocol schemes, or different data capturing times as mentioned above).
+Enhancing the data with additional posts: Another limitation of the CrowdTangle extension was its restriction to retrieving data for up to 500 posts per URL.
+To address this, we retrieved additional data from CrowdTangle for URLs with over 500 posts in the public interaction category.
+While some of the pre- and post-processing steps described above require significant effort, they help ensure an accurate and fair comparison between the public and combined sets in such a way that we can provide conclusive insights into the relative sharing patterns of public vs.
+private posts.
+3. Dataset
+High-level summary: After applying the aforementioned filtering steps, we have in total 19,050 articles from 1,121 news outlets remaining in our dataset.
+These articles have been shared in 253,350 posts, which combined are responsible for 225,282,745 interactions (out of which 70,428,316 are public).
+Bias data: Table 1 shows the number of articles and the total combined and public interactions for each bias and reliability class.
+With the selected bias thresholds, our dataset includes 4,516 Left-biased articles and 3,440 Right-biased articles, with the remaining 11,094 articles falling in the Center category.
+These articles are in turn responsible for 79,038,246 (Left), 56,160,101 (Right), and 90,084,398 (Center) interactions.
+While the Center category includes 58% of the articles, it is interesting to note that it is responsible for a significantly smaller fraction of the total interactions (40%).
+Instead, the Left-biased and Right-biased articles see relatively higher interaction rates (studied in the next section).
+For example, the Left-biased articles account for only 24% of the articles but 35% of the interactions, and the Right-biased articles are responsible for only 18% of the articles but 25% of the interactions.
 Reliability data: For the reliability classes as shown in Table 1, the relative differences are smaller.
 Here, the Reliable articles make up the largest share (9,308 articles and 89,596,137 interactions), followed by Unreliable (3,795 articles and 46,712,927 interactions), Most-unreliable (3,520 articles and 45,299,475 interactions), and Most-reliable (2,427 articles and 43,674,206 interactions).
-4 Public vs. Private Interactions
-To better understand the differences in interactions with public vs.
-private posts, we first compare the interaction rates with all public posts and with all combined (public + private) posts.
-We note that this implicit comparison approach provides more stable and reliable statistical comparisons than trying to use only the estimated number of private interactions.
-The relative strength of this approach is further compounded by the private interactions typically dominating the combined set.
-4.1 Aggregated (Macro) Analysis of Interactions
-Figure 3: Normalized number of combined and public interactions per article for each bias and reliability class.
+4. Public vs. Combined Interactions
+To understand how engagement patterns differ between public and private spheres on Facebook, we present two complementary analyses: (1) an aggregated comparison of public vs.
+combined (public + private) interactions (Sec.
+4.1), and (2) a statistical analysis of their differences (Sec.
+4.2).
+By comparing public interactions against combined interactions, where private interactions constitute the majority (as we show later), we can effectively study how engagement in private spaces differs from public ones.
+4.1. Aggregated (Macro) Analysis of Interactions
 Figure 2 presents the number of combined and public interactions per article for each class. Here, two metrics are used:
-  • Average combined interactions (left bars for each class) represents the average number of combined (private+public) interactions per article in each class.
+Average combined interactions: (left bars for each class) represents the average number of combined (private+public) interactions per article in each class.
 More specifically, considering $I_{c,i}^{\text{comb}}$ as the combined interaction for the $i^{th}$ article in class $c$, the average combined interactions for this class ($\bar{I}_{c}^{\text{comb}}$) is calculated as:
-    $\bar{I}_{c}^{\text{comb}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}{N_{c}},$ (1)
+(1) $\bar{I}_{c}^{\text{comb}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}{N_{c}},$
 where the numerator ($\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}$) corresponds to the “Combined interactions” column in Table 1 and the denominator (i.e., the total number of articles $N_{c}$ for class $c$) is found in the “Articles no” column of Table 1.
 For example, for the Center class in the Bias category, the average combined interactions per article is computed by dividing the total interactions (90,084,398) by the articles count (11,094), yielding $\bar{I}_{\text{Center}}^{\text{comb}}$ to be 8,120 combined interactions per article (gray bar in Figure 2).
-  • Average public interactions per article (right bars) is defined in a similar manner but considering only the public interactions.
+Average public interactions per article: (right bars) is defined in a similar manner but considering only the public interactions.
 More specifically, considering $I_{c,i}^{\text{public}}$ as the public interactions for the $i^{th}$ article in a class $c$, the average public interactions for this class ($\bar{I}_{c}^{\text{public}}$) is calculated as follows:
-    $\bar{I}_{c}^{\text{public}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{N_{c}},$ (2)
+(2) $\bar{I}_{c}^{\text{public}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{N_{c}},$
 where the numerator (i.e., $\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}$) corresponds to the “Public Interactions” column in Table 1.
@@
 the combined statistics.
-First, for all three bias classes and for all four reliability classes, the interaction rates are significantly lower for the public posts than for the combined set (which for all classes are dominated by private posts).
+First, for all three bias classes and for all four reliability classes, the interaction rates are significantly lower for the public posts than for the combined set.
 These aggregate rate differences suggest that users are more likely to interact with private posts, regardless of bias and reliability class, and underscores the importance of considering the total (combined) interactions, not only the (typically studied) public posts.
@@
 combined sets for such interaction comparisons.
-Bias comparisons: Second, looking at the relative differences among the bias classes, the Left bias class has the highest average combined interactions per article (17,501), followed by the Right (16,325) and Center (8,120) classes.
+Bias comparisons: Second, examining the relative differences among the bias classes, the Left bias class has the highest average combined interactions per article (17,501), followed by the Right (16,325) and Center (8,120) classes.
 For public interactions; however, the Right bias class has the highest average public interactions per article (5,236), with the Left (4,497) and Center (2,894) classes trailing behind.
-This shows that biased articles see even greater relative interaction rates when posts are private than public.
+This shows that biased articles see even greater relative interaction rates in private than in public posts.
 Reliability comparisons: Third, for the reliability classes, the Most-reliable class has the highest average combined and public interactions per article (17,995 and 5,101).
@@
 Moreover, the Most-unreliable class has a higher average combined interactions per article (12,869) and public interactions per article (3,751).
-It is interesting to see that the two extreme classes (i.e., the Most-reliable and Most-unreliable) have the highest average interactions per article in both combined and public contexts, implying that users are more likely to engage with articles on both ends of the reliability spectrum, regardless if the posts are public or non-public, although the level of interactions is different.
-Key observation: Compared to the Center group, biased groups receive higher levels of aggregated engagement per article.
-In terms of reliability, the Reliable news reaches the lowest engagement levels per article.
-Aggregated public interaction share: Motivated by most prior works only studying the public posts tracked by CrowdTangle, we next look closer at how big fraction of the total interactions this set captures and compare the relative differences in the public interaction share of the different categories.
-For this analysis, we calculate the (aggregate) public interaction share, defined as the ratio of public interactions to the combined interactions for each class $c$ as $\bar{I}_{c}^{\text{public}}/\bar{I}_{c}^{\text{comb}}$.
-Figure 4: Aggregated public interactions shares
+It is interesting to see that the two extreme classes (i.e., the Most-reliable and Most-unreliable) have the highest average interactions per article in both combined and public contexts, implying that users are more likely to engage with articles on both ends of the reliability spectrum.
+Aggregated public interaction share: Motivated by most prior works only studying the public posts tracked by CrowdTangle, we next evaluate what fraction of the total interactions this set captures and compare the relative differences in the public interaction share of different categories.
+For this analysis, we calculate the (aggregate) public interaction share as the ratio of public interactions to total interactions for each class $c$: $\bar{I}_{c}^{\text{public}}/\bar{I}_{c}^{\text{comb}}$.
 Figure 3 summarizes these ratios.
@@
 However, it should be noted that the two extreme classes (Most-reliable and Most-unreliable) have the lowest (and almost similar) aggregated public interaction shares.
-Key observation: Compared to right-biased articles, left-biased articles receive higher private interactions share.
+Takeaway: Compared to right-biased articles, left-biased articles receive higher private interactions share.
 In terms of reliability, extremely reliable/unreliable news articles receive higher levels of private interactions share.
-4.2 Granular Statistical (Micro) Analysis of Interactions
+Figure 3. Aggregated public interactions shares
+(a) Disributions of different bias classes.
+(b) Disributions of different reliability classes.
+Figure 4.
+Empirical complementary cumulative distribution functions (CCDFs) for combined and public interactions, along with inset boxplots showing the distribution of combined interactions.
+Both CCDFs emphasize tail characteristics of distributions, while the boxplots provide granular statistical summaries, including medians, quartiles, and means.
+4.2. Granular Statistical (Micro) Analysis of Interactions
 Having presented aggregated analysis and insights, we next present a more detailed statistical comparison between different classes in which we give equal consideration to every article in a class, regardless of its total interaction count.
 For this analysis, we utilize the distributions of the per-article level interactions for each class $c$: $\mathcal{D}^{\text{comb}}_{c}=\{I_{c,i}^{\text{comb}}\}$ and $\mathcal{D}^{\text{public}}_{c}=\{I_{c,i}^{\text{public}}\}$, where $I_{c,i}^{\text{comb}}$ is the total number of interactions across all posts associated with the $i^{th}$ article of class $c$ ($1\leq i\leq N_{c}$) and $I_{c,i}^{\text{public}}$ is the corresponding number of interactions with public posts.
-The next two subsections compare the distributions of the total (combined) interactions (Sect.
+The next two subsections compare distributions of the total (combined) interactions (Sect.
 4.2.1) and study the distributions of the relative public interaction shares (Sect.
@@
 In both subsections, the distributions are compared using multiple statistical tests; statistical significance is reported if a p-value is below 0.001, and supporting p-values are reported for the main findings.
-4.2.1 Distribution of Combined Interactions
-[…]
-Figure 5: Empirical complementary cumulative distribution function (CCDF) of both combined and public interactions, alongside an inset boxplot detailing the distribution of the combined interactions for different bias classes.
-While the CCDF underscores the tail characteristics of the distributions, the boxplot provides a granular statistical summary of the combined interactions, including medians, quartiles, and means.
+4.2.1. Distribution of Combined Interactions
+Comparison of Bias Classes: Figure 4(a) shows the CCDFs of both the set of combined interactions ($\mathcal{D}^{\text{comb}}_{c}$) and the set of public interactions ($\mathcal{D}^{\text{public}}_{c}$) for each bias class $c$ on a log-log scale.
+Accompanying the CCDF is a boxplot, highlighting key statistical percentiles for the distribution of the combined interactions ($\mathcal{D}^{\text{comb}}_{c}$).
+Specifically, the 10th percentile (bottom marker), 25th percentile (bottom of box), median (middle marker), 75th percentile (top of box), 90th percentile (top marker), and the mean (i.e., $\bar{I}_{c}^{\text{comb}}$) (circle).
+To allow higher resolution, outliers are not shown in the boxplot.
 Upon examining the figure, we can make several observations.
@@
 Right).
-Key observation: Statistical analysis further strengthens the observation that biased news consistently garners higher combined engagement levels per article than center-aligned news.
+Takeaway: Statistical analysis further strengthens the observation that biased news consistently garners higher combined engagement levels per article than center-aligned news.
 Interestingly, there is no statistically significant difference between the Left and Right class.
-Comparison of Reliability Classes: Figure 6 shows the CCDFs for the different reliability classes, broken down for both combined and public interactions, accompanied by a boxplot offering a clearer perspective on the percentile values for the combined interactions.
+Comparison of Reliability Classes: Figure 4(b) shows the CCDFs for the different reliability classes, broken down for both combined and public interactions, accompanied by a boxplot offering a clearer perspective on the percentile values for the combined interactions.
 The CCDFs again demonstrate heavy-tailed distribution characteristic, and the relative shifts of the distributions are consistent with the aggregate (average values) previously discussed (and seen in Figure 2), with the Most-reliable class typically getting the most interactions, followed by the unreliable classes, and subsequently by the Reliable class.
@@
 For example, here, the core of the boxplot does not distinctly set apart the Most-reliable from the unreliable classes.
-To derive robust conclusions, we therefore again performed a sequence of statistical tests on the combined interactions distributions.
-Figure 6: CCDF of both combined and public interactions, alongside an inset boxplot detailing the distribution of the combined interactions for different reliability classes.
-While the CCDF underscores the tail characteristics of the distributions, the boxplot provides a granular statistical summary of the combined interactions, including medians, quartiles, and means.
+To derive robust conclusions, we, therefore, again performed a sequence of statistical tests on the combined interactions distributions.
 First, the Kruskal-Wallis test was applied to discern differences among the reliability classes.
@@
 Third, the Mann-Whitney U test reinforced our findings, particularly highlighting the statistical significance when comparing the medians of the Reliable class (median of 891) with the others: Most-unreliable (1,162), Unreliable (1,104), and Most-reliable (1,218).
-While all three pairwise comparisons are significant, their p-values range from $8.96\cdot 10^{-9}$ (Reliable vs.
-Unreliable) to $8.4\cdot 10^{-16}$ (Reliable vs.
-Most-reliable).
 Finally, we apply bootstrapping to compare the means.
 While only the Reliable and Most-reliable classes comparison is significant at the 99% confidence level, we note that all three comparisons against the Reliable class are significant at the 95% confidence level.
-Key observation: Comparing to the other reliability classes, the Reliable articles demonstrate statistically significant lower interaction values, especially when compared to the Most-reliable articles.
-4.2.2 Distribution of Public Interaction Shares
+Takeaway: Comparing to the other reliability classes, the Reliable articles demonstrate lower interaction values, especially when compared to the Most-reliable articles.
+4.2.2. Distribution of Public Interaction Shares
 The preceding analysis centered on the distribution of combined interactions.
@@
 For this analysis, we define the set of public interaction ratios (PIR) of class $c$ as $\mathcal{D}^{\text{PIR}}_{c}=\left\{\frac{I_{c,i}^{\text{public}}}{I_{c,i}^{\text{comb}}}\right\}$, where $1\leq i\leq N_{c}$ is the index of articles included in class $c$.
-Figure 7: Distribution of public interaction ratios
-[…]
+Figure 5. Distribution of public interaction ratios
+Figure 5 presents the public interaction ratios using box-plots broken down per bias and reliability category.
+Here, we again show the 10th percentile, 25th percentile, median, 75th percentile, and 90th percentile, as well as the average.
+Like for the aggregate analysis, we observe some notable differences, which we can now support using statistical analysis.
+First, the Kruskal-Wallis test shows that there are statistically significant differences between both the biases classes (p-value of $2.68\cdot 10^{-22}$) and reliability classes (p-value of $8.74\cdot 10^{-33}$).
 Motivated by the higher significance (smaller p-value) for the reliability classes, we consider these differences first.
@@
 It is also worth noting that our observation that the two extreme classes have lower public interaction ratios (higher private interaction ratios) aligns with our findings from the aggregated analysis.
-Key observation: The two extreme classes, Most-reliable and Most-unreliable, exhibit similar private interaction ratios that are statistically higher compared to the other classes.
+Takeaway: The two extreme classes, Most-reliable and Most-unreliable, exhibit similar private interaction ratios that are statistically higher compared to the other classes.
 We now shift our attention to the public interaction ratios across the bias classes, where we make some interesting observations regarding the classes’ relative order.
 Specifically, we now observe the first difference in the relative order of the classes when comparing the average (and median) public shares calculated on a per-article basis (Figure 5) with the aggregated results (Figure 3).
-More specifically, comparing with the aggregate results, the Right class has now switched ranking with the Left class to become the class with the noticeably smallest public interaction share.
+More specifically, comparing with the aggregate results, the Right class has switched ranking with the Left class, becoming the one with smallest public interaction share.
 Furthermore, comparing the distributions seen for the Right and Left class is statistically significant: Dunn test (p-value of $3.25\times 10^{-14}$), Mann-Whitney U test (p-value of $6.16\times 10^{-17}$) and bootstrap results being significant at 99% level.
+Figure 6. Deep interaction ratios (DIRC and DIRP) for different bias classes and a baseline.
+Figure 7. CDFs of DIRC (combined) and DIRP (public) for different bias classes.
+[…]
+Figure 8. CDFs of DIRC (combined) and DIRP (public) for different reliability classes.
+[…]
 While these results may appear surprising at first, they are due to some interesting differences in the distributions.
@@
 Second, taking a closer look at the distributions, $\mathcal{D}_{Left}^{PIR}$ and $\mathcal{D}_{Right}^{PIR}$ differ substantially.
-Part of this can be observed in Figure 5, where we can see that Left class demonstrates notable disparities between public and combined interactions in the tail, while the Right class exhibits relatively larger differences in the head of the distribution.
+Part of this can be observed in Figure 4(a), where we can see that Left class demonstrates notable disparities between public and combined interactions in the tail, while the Right class exhibits relatively larger differences in the head of the distribution.
 This suggests that articles with higher combined number of interactions in the Right class may exhibit higher normalized public interaction shares, and articles with few combined interactions may exhibit relatively smaller public shares (compared to the Left class).
@@
 This confirms that the articles in the Right class with the highest interactions have elevated normalized public interaction ratios, and those in the Left class with the lowest interactions have similarly elevated ratios.
-Key observation: In terms of bias, the Right class sees the smallest public interaction share among the least popular content to interact with, while the Left class sees relatively less overall public interaction share (due to its most popular articles to interact with not seeing as big public interaction share).
-5 Deep vs. Shallow Interactions
+Takeaway: In terms of bias, the Right class sees the smallest public interaction share among the least popular content to interact with, while the Left class sees relatively less overall public interaction share (due to its most popular articles to interact with not seeing as big public interaction share).
+5. Deep vs. Shallow Interactions
 In this section, we investigate the prevalence of deep vs.
 shallow interactions and explore their variations among different reliability and bias classes.
-For this analysis, we categorize all (emoji-based) reactions (i.e., likes, loves, sads, etc.) as shallow interactions while comments and shares are considered as deep interactions.
-This distinction, supported by the previous works [8, 33], is motivated by the idea that comments and shares typically involve more cognitive effort and engagement from users, as they require formulating thoughts, opinions, or responses in the form of comments or actively endorsing and disseminating the content by sharing.
-In contrast, reacting with an emoji is a more passive and less cognitively demanding form of interaction.
+For this analysis, we categorize all (emoji-based) reactions (e.g., likes, loves, sads) as shallow interactions while comments and shares are considered as deep interactions.
+This distinction, supported by the previous works (Aldous et al., 2019; Kim and Yang, 2017), is motivated by the idea that comments and shares typically involve more cognitive effort and engagement from users.
+For example, comments involve formulating thoughts and opinions, while shares actively endorse and disseminate content.
 To compare the depth of interaction seen for subsets of public and private posts, we define the following two per-article metrics:
@@
   • Deep Interactions Ratio for Public (DIRP): For each article, this ratio measures the proportion of the public interactions that are classified as deep interactions.
-For example, consider an article with 1,000 total interactions, 500 of which are deep interactions, resulting in a DIRC of 0.5.
-Now, if 400 of these interactions are public, and 100 of them are deep interactions, the DIRP is 0.25 for the same article.
+To compute DIRC, we divide the number of deep interactions by the total interactions for each URL.
+For DIRP, we first identify the deep interactions (shares and comments) occurring on public posts associated with the URL.
+Summing these deep interactions across all public posts and dividing by the total number of public posts’ interactions for the URL yields the DIRP.
+For example, an article with 1,000 total interactions, 500 of which are deep, has a DIRC of 0.5.
+If 400 interactions are public, and 100 of those are deep, the DIRP is 0.25.
 We next analyze the distribution of DIRC and DIRP values for all articles in each class.
@@
 As a baseline, we also include the distributions for the general population (rightmost box pair), considering all samples.
-Figure 8: Deep interaction ratios (DIRC and DIRP) for different bias classes and a baseline.
-First, referring to the figure, for each bias class, we observe substantial difference between the distributions of DIRC and DIRP.
-For example, for every percentile and for the means, the DRIC distributions (combined) have significantly larger values than the DIRP distributions (public).
+First, referring to the figure, for each bias class, we observe substantial differences between the distributions of DIRC and DIRP.
+For example, for every percentile and for the means, the DIRC distributions (combined) have significantly larger values than the DIRP distributions (public), indicating that users engage in more substantive interactions in less-public settings.
+This suggests that deeper engagement—through comments and shares—tends to comparatively have higher ratios in spaces with fewer visibility constraints, where users may feel more comfortable expressing opinions or endorsing content.
+This pattern is consistent across all bias (Figure 8) and reliability classes (Figure 8), reinforcing the notion that private interactions are an essential component of online engagement dynamics.
 This again highlights the difference between the public and combined interactions and the importance of considering both the private and public spheres in future research.
 This observation leads us to the following key insight.
-Key observation: Irrespective of the news class, users tend to engage more deeply in private discussions compared to public ones.
+Takeaway: Irrespective of the news class, users tend to engage more deeply in private discussions compared to public ones.
 Now, and perhaps even more importantly, to better understand to what extent the relative level of interaction seen for different bias classes depends on which dataset was used we apply the same statistical testing methodology as discussed in previous sections for the two sets independently.
-First, the Kruskal-Wallis test provides evidence that the distributions differ significantly among the bias classes, regardless if using of the DIRC and DIRP, with p-values of $7.14\cdot 10^{-23}$ and $2.08\cdot 10^{-9}$, respectively.
+First, the Kruskal-Wallis test provides evidence that the distributions differ significantly among the bias classes, regardless of using the DIRC and DIRP, with p-values of $7.14\cdot 10^{-23}$ and $2.08\cdot 10^{-9}$, respectively.
 Comparing the p-values, we note that the differences appear more significant using the combined set (DIRC) than using only the public data (DIRP).
-Second, consistent with visual comparison in the figure, the post-hoc Dunn test only supports the statistically significance differences between the Right group (which has the biggest fraction of deep interactions regardless of using the combined or pubic sets) and the other two groups but not among the Center and Left group.
+Second, consistent with visual comparison in the figure, the post-hoc Dunn test only supports statistically significant differences between the Right group (which has the biggest fraction of deep interactions regardless of using the combined or pubic sets) and the other two groups but not among the Center and Left group.
 Finally, the Mann-Whitney U test and bootstrapping support the same pattern for the median and means, with the only significant differences again being between the Right class and the other two bias classes.
 For the bias classes, the main results regarding the relative depth of the interactions among the classes are, therefore, consistent regardless of the dataset.
-Figure 9: CDFs of DIRC (combined) and DIRP (public) for different bias classes.
 To illustrate the tendency of the posts with Right-biased articles seeing a higher deep interaction share, regardless of whether publicly shared or not, we include the CDFs of DIRC (combined) and DIRP (public) for all three biases classes in Figure 8.
 We note that in both cases, the Right class exhibits a noticeable shift to the right, supporting the following insight.
-Key observation: Facebook users tend to engage more deeply with content from the Right bias class in both public and combined contexts.
+Takeaway: Content from the Right bias class in both public and combined contexts gets deeper user interactions.
 Comparison of reliability classes: We now shift our focus to the reliability classes and their deep interaction ratios.
 Figure 8 summarizes these results.
-While we again observe deeper interactions for the combined set (suggesting shallower interactions with public posts), the relative differences among the reliability classes themselves (given a dataset) are less apparent compared to those observed among the bias classes.
+While we again observe deeper interactions for the combined set (suggesting shallower interactions with public posts), the relative differences among the reliability classes themselves are less apparent compared to those observed among the bias classes.
 We also observe some smaller differences in which relative differences are significant, highlighting the value of also considering the combined data (not only the public data).
-Figure 10: Deep interaction ratios (DIRC and DIRP) for different reliability classes
-Figure 11: CDFs of DIRC (combined) and DIRP (public) for different reliability classes.
-First, starting with the combined dataset (DIRC), we observe presence of significant differences among the distributions (Kruskal-Wallis test), with the post-hoc Dunn test (distribution) and Mann-Witney U test (median) revealing that only the Most-unreliable class differs statistically from the others.
-This observation is further visually supported by the CDF plot in Figure 11, which clearly shows a shift to the right in the DIRC distribution of the Most-unreliable class.
-Switching our focus to the DIRP distributions, containing only public interactions, we interestingly observe a slightly different pattern, as we, for that dataset, also observe statistical differences between some additional classes.
+First, starting with the combined dataset (DIRC), we observe the presence of significant differences among the distributions (Kruskal-Wallis test), with the post-hoc Dunn test (distribution) and Mann-Witney U test (median) revealing that only the Most-unreliable class differs statistically from the others.
+This observation is further visually supported in Figure 8, which clearly shows a shift to the right in the DIRC distribution of the Most-unreliable class.
+Switching our focus to the DIRP distributions, containing only public interactions, we interestingly observe a slightly different pattern, as we here also observe statistical differences between some additional classes.
 First, as suggested by the DIRP CDFs in Figure 8, both extreme classes (Most-unreliable and Most-reliable) are statistically different from the other two classes (Kruskal-Wallis test followed by Mann-Whitney test, followed by the post-hoc Dunn test).
@@
 We base the following insight based on our observations across the two datasets.
-Key observation: Users exhibit higher deep interaction levels with content from the Most-unreliable class.
-6 Correlation-based Analysis
-Given the significant differences observed in interaction rates and deep interaction levels between public and private posts, we aim to determine if these variations stem from specific articles within each bias/reliability class that receive more or fewer private interactions or if certain articles consistently exhibit higher (deep) interaction rates, regardless of public or private sharing.
-To investigate this, we next examine the correlations between public post statistics and their combined counterparts.
-Public vs.
-combined interactions: Table 3 summarizes the correlation statistics (between combined and public interactions) observed for each class using two different correlation metrics: Spearman and Kendall.
-Given that the normality assumption required for the Pearson correlation is not satisfied for the different classes, we opted for non-parametric correlation metrics.
-These metrics are better suited to capture monotonic relationships, account for outliers, and take into consideration the ranks of the data points.
-Table 3: Correlation between public and combined interactions. (Bold/italics indicate highest/lowest correlation.)
-Class Spearman Kendall
-Bias Left 0.82 0.66
-Center 0.74 0.59
-Right 0.74 0.58
-Reliability Most-unreliable 0.73 0.58
-Unreliable 0.79 0.63
-Reliable 0.76 0.61
-Most-reliable 0.77 0.62
-While different trends are observed for reliability classes, a consistent pattern emerges for the bias classes across both correlation metrics.
-The Left class exhibits the strongest correlations, followed by the Center class, and finally, the Right class with the weakest correlations.
-This suggests that articles with a left-leaning bias tend to have more similar sharing and interaction patterns in both public and private contexts, indicating higher similarity in engagement across different articles.
-In contrast, articles with a right-leaning bias show more variability in their public and private interactions, which may reflect a more diverse range of opinions or reactions to the content among readers.
-These observations are consistent with our observations in Section 4.2.2, which showed that the Right class saw much more diverse public interaction shares depending on the level of combined interactions of the article (see our bucket-based analysis).
-Table 4: Correlations for deep interaction shares: public vs combined. (Bold/italics indicate highest/lowest correlation.)
-Class Spearman Kendall
-Bias Left 0.64 0.48
-Center 0.62 0.47
-Right 0.51 0.39
-Reliability Most-unreliable 0.55 0.42
-Unreliable 0.61 0.47
-Reliable 0.64 0.49
-Most-reliable 0.57 0.42
-Deep interaction rates: We next consider the correlation between the deep interactions ratio for combined (DIRC) and public (DIRP) posts. Table 4 summarizes these results.
-First, when comparing the correlations for deep interaction ratios across the bias classes, we observe that the Left and Center bias classes show higher correlations across both metrics, while the Right class exhibits weaker relationships.
-This pattern suggests that deep interactions, or more in-depth engagements with content, are more consistent between public and private contexts for left-leaning and centrist articles.
-Second, in terms of reliability, the correlations are relatively similar across classes, with the Reliable class showing the strongest relationships and the Most-unreliable class having the weakest.
-These findings indicate that the consistency of deep interactions across public and private contexts is less affected by articles’ perceived reliability than their political bias.
-7 Discussion
-Our research provides a comprehensive exploration of news interactions on Facebook, with a specific focus on discerning the nuances between public and combined interactions across diverse news classes.
-We developed a robust comparison methodology, crucial for future studies in this area.
-Our findings, presented in several key facets, provide valuable insights along several dimensions (discussed next).
-Reliability and Bias: Several discoveries emerged when examining the reliability of news articles.
-Surprisingly, articles classified as Reliable garnered lower levels of engagement compared to those classified as Most-reliable or even various Unreliable categories.
-Furthermore, our investigation into the impact of biases unveiled that biased news, regardless of whether it leans Left or Right, consistently received higher levels of user interaction than more neutral (Center) content.
-Public vs.
-Private: One of the central aspects of our study concerned the comparisons of the interaction patterns within the public vs.
-private spheres.
-What became evident was that articles classified as Most-reliable or Most-unreliable, representing the extremities of the reliability spectrum, attracted significantly higher proportions of engagement within private spaces compared to their public counterparts.
-Intriguingly, no statistically significant differences were observed between these two extreme groups here.
-This observation holds substantial implications for content moderators, policymakers, and researchers, prompting us to ponder whether an exclusive focus on the public facet of Facebook provides a comprehensive understanding of user interactions, particularly concerning news content classified as Most-unreliable.
-Interaction Depth: Our research extended beyond interaction frequency to explore the depth of interactions, revealing noteworthy variations between the public and private spheres.
-Notably, statistically significant differences in interaction depth were observed, reinforcing the importance of considering both spheres in research endeavors.
+Takeaway: Users exhibit higher deep interaction levels with content from the Most-unreliable class.
+6. Topic-Based Analysis
+In this section, we extend our investigation by conducting a topic-based analysis of news articles to assess whether engagement patterns vary across different content themes.
+For this analysis, we first extracted the textual content from the dataset news articles, a process that succeeded for 14,996 items.
+We then employed the ChatGPT 4o-mini model to classify each article into one of 14 predefined topics: Politics, Health & Medicine, Crime, Business  & Finance, Environment &  Climate, Entertainment, Education, Sports, Science, Lifestyle & Leisure, Religion, Technology, Arts &  Culture, and Food.
+This specific set of topics was selected to capture the diverse range of subject matters prevalent in contemporary news consumption, while also reflecting societal interests and trends.
+Notably, the ordering of topics here mirrors the distribution observed in our dataset—ranging from the most frequent (Politics, with 7,956 articles) to the least frequent (Food, with 57 articles).
+However, it is important to note that these frequencies are influenced by two key factors: (1) the initial article selection criteria of Ad Fontes Media, and (2) the timeframe of our dataset, which spans a period heavily impacted by the COVID-19 pandemic, leading to an elevated presence of Health & Medicine content (the second most prevalent topic).
+Finally note that for articles where none of these predefined topics were clearly identifiable, the model was instructed to leave them uncategorized.
+In total, 14,650 articles (97.7%) were successfully classified into one of the 14 categories.
+First, to assess whether our earlier findings hold within specific topics, we conducted a stratified analysis across topic groups.
+Our results indicate that the previously observed patterns remain, with statistical significance for several of the classes with enough samples.
+For example, in the case of the Politics topic the insights reported in both Sects.
+4 and 5 remain statistically robust when considering the reliability dimension.
+As another example, for Health & Medicine (the 2nd largest topic), all the conclusions in Sects.
+4 and 5 for both the bias and reliability dimensions are still statistically significant.
+These confirm the general validity of our broader findings within many of these topics.
+Second, to explore how engagement varies across topics, we analyzed the distributions of the public interaction ratio for each topic.
+This analysis was conducted independent of the articles’ bias and reliability, as some topics had too few samples to allow for a breakdown by those categories.
+Figure 9 shows the distribution of the public interaction ratios (as a boxplot) for each of the topics, sorted based on their medians.
+From this figure, we can first see that Entertainment-related articles exhibit the highest public interaction ratio, indicating that content from this category is more likely to be engaged with in public discussions.
+This suggests that entertainment news, which often includes celebrity updates and viral stories, is inherently more shareable and widely disseminated in public spaces.
+On the other hand, Technology-related articles exhibit the lowest public interaction share, meaning they receive a disproportionately high share of engagement in private spaces.
+This may be attributed to the technical nature of these articles, which are often discussed within niche communities or shared in professional or interest-based private groups.
+Figure 9. Boxplots of the public interaction ratio for different topics (sorted by median).
+Figure 10. Heatmap of Kruskal-Wallis test results for topics’ public interaction ratios.
+It is important to note that not all observed differences in public interaction ratios between topics in Figure 9 are statistically significant.
+To validate the statistical significance of these observed differences, we conducted pairwise Kruskal-Wallis tests between topic groups.
+Figure 10 presents a heatmap illustrating the statistical significance of differences in public interaction ratios across topics.
+Orange-colored cells here indicate significant differences ($p<0.1$), while gray cells indicate non-significant comparisons.
+This analysis confirms that Entertainment-related content has a significantly higher public interaction share compared to all other topics, whereas Technology-related content exhibits significantly lower public engagement compared to most categories.
+Conversely, topics such as Food, we could not find statistically significant differences in public interaction ratios compared to other topics (except for Entertainment).
+It is worth noting that we selected a p-value threshold of 0.1 for our analysis, as stricter thresholds (e.g., 0.01) did not yield many statistically significant differences between topics.
+This more lenient threshold was chosen primarily to compensate for the limited sample size for certain topics in our dataset.
+7. Top Publishers Analysis
+To further contextualize our findings, we analyze the public interaction ratio (PIR) and the deep interaction ratio (DIRC) for the top publishers in our dataset.
+In this context, “top publishers” refers to those with the highest number of articles in our (and Ad Fontes Media) dataset.
+For this analysis, we excluded Yahoo, as it primarily functions as a news aggregator and does not produce original content.
+We begin by examining the public interaction ratio across the top publishers.
+Figure 11(a) presents a boxplot of the public interaction ratio for each outlet, sorted by median values.
+For easier interpretation, the color of each “box” indicates the bias class (i.e., Left, Center, or Right) associated with each publisher (based on the labeling of the articles they have published).
+It is worth noting that some outlets, such as Politico and NBC News, may shift between the Left and Center classifications over time due to editorial changes or evolving media landscapes.
+From this figure, we observe that New York Post exhibits the lowest public interaction ratio, suggesting that a larger proportion of its engagement occurs in private spaces.
+This could be attributed to the nature of its audience, which may prefer engaging with its content in more private settings, such as closed groups.
+Additionally, the sensationalist reporting of the New York Post might encourage discussions that users feel more comfortable having in less-visible online spaces.
+Next, we investigate which of these publishers generate the deepest user engagement, as measured by the deep interaction ratio (DIRC).
+Figure 11(b) illustrates the DIRC distributions for the top publishers, again sorted by median values.
+Notably, New York Post and Fox News, both from the Right class, exhibit the highest median DIRC values, followed by CNN and New York Times from the Left class.
+At the lower end of the spectrum, we find NPR and Reuters, both from the Center class.
+In general, we observe that publishers with the highest DIRC values tend to be biased ones, while those from the Center class have the lowest values.
+The probability of this pattern occurring if we randomly sort the 10 publishers is approximately 2.22%.
+This observation may suggest that users are more likely to engage deeply with content from publishers that have a clear political leaning, compared to those that maintain a more neutral stance.
+(a) Boxplots of the public interaction ratios (sorted by median).
+(b) Boxplots of the deep interaction ratios (sorted by median).
+Figure 11. Comparison of (a) public interaction ratio and (b) deep interaction ratio for the top-10 outlets.
+8. Related Works
+This research most closely relates to studies exploring user engagement dynamics with various social media content.
+Our main contribution lies in the novel analysis of interaction disparities between public and private news article sharing on Facebook, particularly concerning articles with varying bias and reliability.
+Examining interactions depth adds another layer to our analysis and findings.
+There exist studies that partially address some dimensions highlighted in this paper.
+However, as discussed in the introduction, the vast majority of these works, particularly pertaining to Facebook, predominantly concentrate on the platform’s public sphere.
+To contextualize our study within the context of the existing literature and to highlight the novelty of our contribution, we next describe the most closely related works.
+One line of this prior research has studied the differences in the engagement levels of different bias or reliability (or both) classes on different platforms including Facebook (Edelson et al., 2021; Hiaeshutter-Rice and Weeks, 2021; Altay et al., 2022), Reddit (Weld et al., 2021; Bond and Garrett, 2023), and Twitter(Mohammadinodooshan and Carlsson, 2024; Vosoughi et al., 2018).
+For example, Edelson et al.
+(Edelson et al., 2021), while focusing on the public sphere, conducted a large-scale study on user engagement with 7.5 million posts from 2,551 publishers (Facebook pages) across both bias and reliability dimensions.
+Their results show that individual posts from non-misinformation news outlets tend to attract lower median engagement than misinformation, aligning with our findings.
+Weld et al.
+(Weld et al., 2021) carried out a similar investigation on Reddit rather than on Facebook.
+When considering the reliability parameter, Vosoughi et al.
+(Vosoughi et al., 2018) analyzed a dataset of approximately 126,000 stories tweeted by over 3 million people more than 4.5 million times, finding that false news stories reached more people than true news and diffused significantly farther, faster, deeper, and more broadly than the truth across various categories of information.
+Notably, during the critical period of the COVID-19 pandemic and by doing a cross-national study, Altay et al.
+(Altay et al., 2022) delineated the surge in online news consumption, with credible news outlets witnessing a significant boost.
+Samory et al.
+(Samory et al., 2020) characterized the social media news sphere through user co-sharing practices by focusing on 639 news sources, both credible and questionable, and divided them into 4 clusters, and characterizing them according to the audience that shares their articles on Twitter and how the stylometric features used by each cluster is successful in getting users engagement.
+Lamot et al.
+(Lamot et al., 2022) examined how news headlines are remediated on Facebook and how this affects user engagement.
+Finally, Boukes et al.
+(Boukes et al., 2022) compared user-content interactivity and audience diversity across news and satire, finding differences in online engagement between satire, regular news, and partisan news.
+Considering the bias dimension, on the other hand, Wischnewski et al.
+(Wischnewski et al., 2021) discovered that users are more inclined to share hyperpartisan news articles that coincide with their own political views.
+The bias towards the right party in sharing and engagement with news is another thesis that has been studied by previous works on both Twitter (González-Bailón et al., 2022) and Facebook (González-Bailón et al., 2023).
+Research has also examined other dimensions of user engagement, with, for example, Aldous et al.
+(Aldous et al., 2019) showing that content topics influence engagement.
+Studies have found that emotional content generates higher Facebook engagement, with Maier et al.
+(Maier, 2015) demonstrating this effect while controlling for author influence.
+Finally, some works have considered the time factor and analyzed temporal patterns of engagement with different news classes (Vassio et al., 2022).
+In summary, while previous research has partly explored some dimensions of user engagement with news, our study is the first to directly compare the dynamics of public and private sharing of news articles on Facebook.
+9. Conclusions
+In conclusion, this study makes several important contributions to the understanding of news engagement on social media.
+By developing a robust methodology through collaboration with the CrowdTangle team, we present the first comprehensive comparison of engagement dynamics between Facebook’s publicly tracked content and overall platform interactions.
+Our analysis of over 19K news articles reveals that public engagement patterns often fail to reflect platform-wide behavior, with significant implications for research methodology and content moderation strategies.
+Our findings challenge several assumptions about news engagement on Facebook.
+First, users consistently engage more deeply with content in less public spaces, regardless of the content’s bias or reliability classification.
+Second, particular attention should be paid to right-biased and least reliable news content, which generate notably deeper engagement across both public and less public spheres.
+Third, we highlight some important limitations of focusing solely on Facebook’s public sphere.
 For example, the extremities of the reliability spectrum (Most-reliable and Most-unreliable) exhibited a tendency to elicit deeper interactions from users within the public sphere.
 However, this narrative shifted when examining combined interactions, indicating that only the Most-unreliable class triggered statistically significant deeper engagement from Facebook users.
-This underscores the intricate interplay between reliability, bias, and interaction depth, urging researchers to explore these dynamics beyond the confines of public engagement.
-Ethical Considerations: It is essential to acknowledge the ethical dimension of our research.
-While our findings offer invaluable insights, they also hold the potential for misuse.
-To mitigate this risk, we undertook several precautions, such as selecting a diverse set of articles and news sources and maintaining objectivity in our analysis.
-We adhered to relevant privacy regulations, ensuring that no personally identifiable information was collected or disclosed.
-Transparency is of utmost importance to us, and we will provide all codes used in this paper, along with the list of news articles, in a camera-ready version.
-Although we cannot publish raw engagement data due to CrowdTangle terms of service, we will furnish attachment status information for URLs and posts through a lookup table with hashed keys.
-This approach enables responsible utilization of our research findings while minimizing the likelihood of negative consequences.
-Beneficiaries: Our findings hold value for a diverse range of beneficiaries.
-Researchers, journalists, and policymakers can gain insights into news engagement patterns, aiding in informed decision-making.
-Social media content moderators can benefit by understanding where to focus their efforts.
-Moreover, our insights have relevance for platforms and media organizations, guiding them in shaping content strategies for more effective audience engagement.
-In conclusion, our research sheds light on complex engagement dynamics on social media and emphasizes the responsible use of these findings for informed decision-making.
-8 Related Works
-This research most closely relates to studies exploring user engagement dynamics with various social media content.
-Our main contribution lies in the novel analysis of interaction disparities between public and private news article sharing on Facebook, particularly concerning articles with varying bias and reliability.
-Examining interactions depth provides another level of depth to our analysis and findings.
-There exist studies that partially touch upon some dimensions highlighted in this paper.
-However, as discussed in the introduction, the vast majority of these works, particularly pertaining to Facebook, predominantly concentrate on the platform’s public sphere, neglecting to provide a comparative analysis with the private or combined sphere.
-To contextualize our study within the context of the existing literature and to highlight the novelty of our contribution, we next describe the most closely related works.
-One line of this prior research investigates user engagement with different classes of news, particularly regarding bias and reliability.
-For example, Edelson et al.
-[7], while focusing on the public sphere, conducted a large-scale study on user engagement with 7.5 million posts from 2551 publishers (Facebook pages) across both bias and reliability dimensions.
-Their results show that individual posts from non-misinformation news outlets tend to attract lower median engagement than misinformation.
-Weld et al.
-[34] carried out a similar investigation on Reddit rather than on Facebook.
-When considering the reliability parameter, Vosoughi et al.
-[35] analyzed a dataset of approximately 126,000 stories tweeted by over 3 million people more than 4.5 million times, finding that false news stories reached more people than true news and diffused significantly farther, faster, deeper, and more broadly than the truth across various categories of information.
-Notably, during the critical period of the COVID-19 pandemic and by doing a cross-national study, Altay et al.
-[36] delineated the surge in online news consumption, with credible news outlets witnessing a significant boost.
-Samory et al.
-[37] characterized the social media news sphere through user co-sharing practices by focusing on 639 news sources, both credible and questionable, and divided them into 4 clusters, and characterizing them according to the audience that shares their articles on Twitter and how the stylometric features used by each cluster is successful in getting users engagement.
-Lamot et al.
-[38] examined how news headlines are remediated on Facebook and how this affects user engagement.
-Finally, Boukes et al.
-[39] compared user-content interactivity and audience diversity across news and satire, finding differences in online engagement between satire, regular news, and partisan news.
-Considering the bias dimension, on the other hand, Wischnewski et al.
-[40] discovered that users are more inclined to share hyperpartisan news articles that coincide with their own political views.
-The bias towards the right party in sharing and engagement with news is another thesis that has been studied by previous works on both Twitter [41] and Facebook [42].
-Other dimensions of user engagement have also been considered.
-For example, Aldous et al.
-[8] investigated the effects of topic and emotional factors on posting across five social media platforms (Facebook, Instagram, Twitter, YouTube, and Reddit).
-They found that user engagement is strongly influenced by the content’s topic, with certain topics being more engaging on a particular platform.
-By controlling the effect of the author (considering articles by the same author) over a one-year period, Maier et al.
-[43], among other things, found that highly emotional topics get a higher level of Facebook engagement.
-A large body of research has also focused on studying the role of sad and anger in social media news sharing and engagement (e.g., [44, 45]).
-Finally, we note that profiling of the relation between user engagement dynamics and different content itself has paved the way for new applications.
-Examples include techniques and models that rely (partly) on user engagement dynamics (as features) for the detection of different content types.
-Such applications may benefit social media content moderators [46, 47].
-Others have profiled the temporal dynamics of user engagement with different classes of news [12, 48].
-In summary, while previous research has partly explored some dimensions of user engagement with news articles, our study is the first to directly compare the dynamics of public and private sharing of news articles on Facebook, specifically examining the interactions and depth of engagement with articles of different levels of bias and reliability.
-9 Conclusions
-In this paper, we presented the first investigation of the differences in interaction patterns between public and private sharing of news articles on Facebook, focusing on articles with varying bias and reliability, as well as the depth of interactions.
-By analyzing a dataset of over 19K news articles (with more than 225M interactions on Facebook), employing the CrowdTangle browser addon for data collection, and carefully processing this data, we provided a comprehensive analysis of user interaction with news content on Facebook.
-Our findings reveal significant differences in interaction patterns between public and private sharing across different classes of news.
-Notably, users tend to engage more deeply in private discussions compared to public ones, irrespective of the bias or reliability class of the news.
-We also found that Facebook users engage more deeply with the least reliable news and the right-biased news, both classes exhibiting higher deep interaction ratio levels.
-While this tendency may suggest that the content from these classes may be more polarizing or controversial, leading to deeper discussions and debates, further research (e.g., by interviewing users) is needed to determine the underlying factors driving this difference in engagement.
-Our research findings offer valuable insights with significant implications for various stakeholders (e.g., researchers, journalists, policymakers, and social media content moderators) and deepen our understanding of how public and private spheres influence the dissemination and engagement with news articles on social media, enabling more informed decision-making.
-For example, our identification of content classes that generate deeper user engagement can provide actionable guidance to Facebook moderators, enabling them to focus their efforts effectively.
-Furthermore, the disparities we uncovered between public and private interaction patterns can guide future research that considers both spheres, promoting a more comprehensive approach.
-To facilitate such analysis, we present a robust methodology developed through active collaboration with CrowdTangle, which ensures a fair comparison between public and private spheres.
-This methodology equips researchers and content moderators with a valuable tool to navigate the evolving landscape of online news engagement.
-In conclusion, by examining the differences in public and private sharing of news articles, this study offers valuable insights into the user engagement patterns associated with contents belonging to different bias and reliability classes, as well as highlights the importance of understanding the dynamics of news sharing and interaction both in the public and private spheres of social media platforms such as Facebook.
-While challenges remain, our methodology and discoveries offer valuable data-driven guidance toward addressing both spheres.
-10 Statements and Declarations
-10.1 Conflict of interest
-The authors have no competing interests.
-10.2 Author contributions
-Both authors have made substantial contributions to the conceptualization, design, data collection, analysis, and writing of the research.
-10.3 Acknowledgment
-The authors thank the CrowdTangle team for their active feedback and help during this research.
-10.4 Data and code availability
-All the codes associated with this research will be made available with the camera-ready version.
-At the same time, to facilitate replicating the results and conducting similar research, the list of all news articles considered in this study and also the attachment status of all related posts will be published.
+This finding suggests that research conclusions drawn solely from public data may not accurately represent overall platform engagement patterns, as not only the magnitude of engagement differs between public and private but also the relative ordering of how different content types engage users.
+These insights carry significant implications for multiple stakeholders.
+For researchers, our findings underscore the importance of considering engagement disparities between public and less public spheres when studying social media behavior, particularly for news content at different ends of the bias and reliability spectrums.
+For content moderators, our results suggest the need to reevaluate strategies based solely on publicly visible interactions.
+For policymakers, our findings highlight the importance of considering both public and less public spheres when developing policies for social media governance.
+Future research should investigate the underlying factors driving these engagement patterns, particularly through user interviews, behavioral studies, and analysis of post comments.
+While challenges remain in understanding social media dynamics, our findings illuminate significant differences in how users interact with news across Facebook’s diverse spheres, highlighting the complexity of news-sharing behavior on social media platforms.
+Acknowledgements.
+The authors express their gratitude to CrowdTangle for providing the Facebook data.
+They also extend their thanks to the anonymous reviewers for their insightful comments that helped improve the paper.
+This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.